### v1

# Comparing Human-Centered Language Modeling: Is it Better to Model Groups, Individual Traits, or Both?

Abstract
Natural language processing has made progress in incorporating human context into its models, but whether it is more effective to use group-wise attributes (e.g., over-45-year-olds) or model individuals remains open. Group attributes are technically easier but coarse: not all 45-year-olds write the same way. In contrast, modeling individuals captures the complexity of each person’s identity. It allows for a more personalized representation, but we may have to model an infinite number of users and require data that may be impossible to get. We compare modeling human context via group attributes, individual users, and combined approaches. Combining group and individual features significantly benefits user-level regression tasks like age estimation or personality assessment from a user’s documents. Modeling individual users significantly improves the performance of single document-level classification tasks like stance and topic detection. We also find that individual-user modeling does well even without user’s historical data.
Comparing Human-Centered Language Modeling:
Is it Better to Model Groups, Individual Traits, or Both?
[…]
1 Introduction
Language varies between people. Two strands of human-centered NLP work have modeled the humans behind the language. The first focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group. These attributes include socio-demographics like age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more. The second strand focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
While these two strands have advanced human-centered NLP (Soni et al., 2023), their relative strengths, complementarity, and impact over different tasks are poorly understood. People are not defined by their group membership alone Orlikowski et al. (2023), but individual traits might not generalize. This paper compares the two approaches and their combination in the same framework and tasks in pre-trained large language models. We use the monolingual socio-demographically adapted model from Hung et al. (2023) and the HaRT model from Soni et al. (2022) for the first two types. We use a multi-task learning setup to create GRIT, a PLM based on HaRT, trained with both individual and group human context in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness). We test all systems on five user- and document-level tasks. Note that because we focus on conceptually comparing group and individual traits, we cannot compare to GPT4, which does not support stratification to either attribute.
PLMs trained on individuals and groups enhance user-level regression tasks like age estimation and personality assessment from user’s multiple documents. Such user-level tasks focus on individual person, and our findings show that these people are best modeled as a combination of their groups and individual traits conforming with the notion that a person is a mix of their group attributes and unique characteristics. Document-level categorization tasks, like stance detection, are more personal. A PLM taught within an individual human context alone improves our considered document-level classification tasks.
By their very nature, models of this kind touch upon sensitive user information. For this reason, we take a responsible release strategy, making only the code for the comparisons publicly available and the exact splits of the TrustPilot and Stance datasets used. We build on top of the publicly available code for HaRT and Hung et al. (2023). We acquired the model and data in a secure manner from the authors of Soni et al. (2022) and TrustPilot data splits from the authors of Hung et al. (2023). For more information about the model and data, see Sections 3 and 4. For a discussion of the ethical implications of the model and data, see Section Ethical Considerations.
Contributions.
Our contributions are: (1) We provide an analysis with a comparison of modeling individual users, group socio-demographic features, and both group and individual traits in PLMs. (2) We evaluate the three modeling strategoies on five downstream tasks: two user-level (personality-openness evaluation and age estimation) and three document-level classification tasks (stance detection, topic detection, and age category prediction). (3) We find that user-level regression tasks like estimating age and assessing personality from user’s multiple documents perform better with mixed individual and group human context (GRIT) and document-level tasks like stance detection and topic detection perform better with individual context alone (HaRT).
2 Integrating Human Context in PLMs
For our comparative study, we use three systems representing the paradigms of human context modeling to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts.
Training with group context.
We build on Hung et al. (2023)’s work to explore demographic adaptation in transformer-based PLMs. They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author. They use the multilingual reviews dataset with demographic labels from Trustpilot [^1] https://www.trustpilot.com/ Hovy (2015). They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection. In our study, we use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4). We use the results from Hung et al. (2023) with the monolingual PLM model BERT on a US-based English dataset with out-of-domain demographic (age) specialization for our comparison study. Out-of-domain data is the Blogs authorship corpus Schler et al. (2006), and in-domain means Trustpilot corpus. To be consistent and fair in comparing with other human context training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
Training with individual human context.
Soni et al. (2022) introduced human language modeling (HuLM), i.e., training a regular language modeling task but including a dynamic individual human context vector derived from the authors’ texts. This vector captures the human states in which the text is generated to induce coherence between different texts generated by the same author. Soni et al. also view this vector as representing the text-derived human factors. They introduce a Human-aware Recurrent Transformer (HaRT), a unidirectional autoregressive PLM that trains for the HuLM task. They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks. We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
Training with both group and individual human context.
We train a PLM that can integrate the author’s individual and group human context knowledge when training for language modeling. We extend Soni et al.’s HaRT model by training for HuLM and predicting a human group attribute in a dynamic muti-task learning setup as used by Hung et al.. We want to induce the individual human context through the author’s language and inject the group context by predicting a group attribute of the author. Predicting group attributes during training can also be viewed as a regularizer for the model, as it constrains the possible output. We discuss the model in detail in section 3. We compare two user-level tasks and a document-level task, as in Soni et al. (2022), and on topic detection and age prediction tasks, as in Hung et al. (2023).
3 Models
This section describes the models we compare to represent group and individual human contexts.
3.1 Modeling individual human context
HaRT.
Soni et al. (2022) introduced HaRT to incorporate individual human context into PLMs. They use a 12-layered autoregressive GPT-2 based architecture with a modified self-attention computation at layer 2. This modification to the query vector now includes the individual human context via a dynamic user-state vector.
[…]
where $IN$ is the insert layer (layer 2), $Q_{i}$ is the query vector under computation, $H_{i}$ is the hidden states vector, and $U_{i-1}$ is the user-state vector derived from the previous block of language seen from the user. All the text from a user is processed in the same forward pass with recurrent processing of blocks of fixed-length (1024) tokens chunked after temporally ordering the social media posts by created time. The user state is recurrently updated using the hidden states from layer 11 and computed as follows:
$U_{i}=tanh(W_{U}U_{i-1}+W_{H}H^{(E)})$
where, $E$ is the extract layer (layer 11), $U_{i}$ is the updated user-state vector, $U_{i-1}$ is the user-state vector from the previous block, and $H^{E}$ is the hidden states vector from layer 11. This formulation of updating the user-state vector extends the previous user-state vector information with the current language block’s information.
HuLM pre-training task.
HaRT is pre-trained for the human language modeling (HuLM) task defined as predicting the next token given the previous tokens while conditioning on previous user state $U_{1:t-1}$ (Soni et al., 2022) .
$Pr(\mathbf{W}_{t}|\mathbf{U}_{t-1})=\prod_{i=1}^{n}Pr(w_{t,i}|w_{t,1:i-1},\mathbf{U}_{1:t-1})$
This is translated into a pre-training objective to maximize:
$\prod_{a\in\textup{Users}}\prod_{t=1}^{|\mathcal{B}_{a}|}\prod_{i=1}^{|B_{t}^{(a)}|}Pr(w_{t,i}|w_{t,1:i-1},B_{1:t-1}^{(a)})$
where, $w_{t,i}$ is the $i^{th}$ token in the $t^{th}$ block ($B_{t}^{(a)}$) for user $a$. The tokens from the previous blocks are represented using HaRT’s recurrently updated user-state vector.
3.2 Modeling group human context
BERTDS and BERTage-MLM.
Hung et al. (2023) explore socio-demographic adapted BERT models to inject group human context into PLMs. We use the names BERTDS and BERTage-MLM to denote their demographic (age) specialization using the multi-task learning setup and demographic adaptation with masked language modeling respectively.
Multi-Task Learning.
Hung et al. (2023) train for domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, they adopt a dynamic MTL objective for training with group human context. Kendall et al. interpret homoscedastic uncertainty as task-dependent weighting and derive a multi-task loss function that can optimally learn the weights to balance the impact of multiple loss functions. This approach accounts for the different scales across multiple losses.
$\tilde{L_{t}}=\frac{1}{2\sigma_{mse}^{2}}L_{t}+\log\sigma_{t}$
Hung et al. minimize the sum of both the uncertainty adjusted losses: $\tilde{L_{mlm}}+\tilde{L_{dem}}$.
3.3 Modeling both individual and group human context
GRIT.
GRIT incorporates both individual and group human contexts using a multi-task learning setup with HaRT that also predicts a socio-demographic attribute of the author in each forward pass using the average of user-state vectors from each non-padded block of the user’s temporally ordered text.
Multi-Task Learning.
GRIT is pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup. The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
$Pr(attribute|\overline{\mathbf{U}})$
Because of compute limitations, we chunk a user’s language history into blocks and process them in a single forward pass. Each block of text from a user results in a user-state vector. We use the average of the user-state vectors from each non-padded block of texts from an author to compute their final user-state representation. This representation is layer-normed and linearly transformed before making a continuous-valued prediction for the specific attribute.
We train one model for the attribute age (GRITage) and one for the attribute personality type openness (GRITope), respectively. The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$). To account for the homoscedastic uncertainty Kendall et al. (2018) of losses, we adopt a dynamic MTL objective as Hung et al. (2023). We use the joint loss for a continuous and discrete output as derived in Kendall et al. (2018) and compute our joint objective as follows:
$\frac{1}{\sigma_{ce}^{2}}L_{ce}+\frac{1}{2\sigma_{mse}^{2}}L_{mse}+\log\sigma_{ce}+\log\sigma_{mse}$
where, $\sigma_{ce}$ and $\sigma_{mse}$ are the variances of the task-specific losses.
In practice, we use log variance for numerical stability and use the adjusted loss calculation as follows:
$\exp^{-\eta_{ce}}L_{ce}+\eta_{ce}+\frac{1}{2}(\exp^{-\eta_{mse}}L_{mse}+\eta_{mse})$
where $\eta_{x}=\log\sigma_{x}^{2}$. We let $\sigma_{ce}$ and $\sigma_{mse}$ be learnable parameters for the model. We do not halve the log-term of the cross-entropy loss since we found it to perform better with our multi-task learning experiments.
Pre-training data.
We use a subset of the pre-training data for HaRT, consisting of the demographics and personality information. This subset contains the Facebook posts from Park et al. (2015) as used by Soni et al.. Our dataset is consistent with the inclusion criteria for HaRT to ensure moderate language history for each user: we include English posts from users with at least 50 total posts and at least 1000 words. This dataset consists of just over 63,000 unique users, which we split into a training dataset consisting of messages from 56,930 users, a development dataset that consists of messages from 1836 users that were not part of the training set, and a test set of messages from a separate set of 4438 users that are neither in training nor the development set. To evaluate the human attribute prediction in GRITope, we use a subset of the test set consisting of messages from 1745 users to accommodate for questionnaire reliability. We use the Facebook posts for the HuLM task and the demographic and personality scores of consenting Facebook users (Kosinski et al., 2013) for the human attribute prediction task.
Training.
We use HaRT’s pre-trained weights as the base weights for GRIT and randomly initialize the newly introduced weights for human attribute prediction. GRIT is trained on our pre-training dataset using the 5e-5 learning rate after experimenting with a few learning rates, including that used for HaRT’s pre-training. Following HaRT, and due to computing limitations, each training instance is capped to 8 blocks of 1024 tokens each, with train batch size as 1 per device and evaluation batch size as 20 per device, trained over 2 GPUs for eight epochs. We explored multiple joint losses before resorting to the homoscedastic loss computation. Since HaRT caps to 4 train blocks for user-level downstream tasks, we also pre-train GRITage and GRITope with four training blocks.
3.4 Transfer Learning
To assess the efficiency of GRIT to transfer learning from predicting one group human attribute to another, we experiment with continuing pre-training for each group attribute. To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness). We do the same for GRITope and continued MTL with predicting age.
3.5 Fine-Tuning
We fine-tune GRIT and HaRT for downstream document-level tasks. Each downstream task has a separate fine-tuned model that is initialized with the respective model’s pre-trained parameters and trained using the respective downstream task labels and an appropriate loss function. We also use the historical language of a user where available for any of the downstream tasks. We use the last predicted token’s representation to predict the label in document-level classification tasks. We experimented with fine-tuning GRIT for user-level regression tasks in multiple ways, including 1) similar to HaRT, by using the averaged user-state vectors from GRIT, 2) same as previous but fine-tuning only the history module, attribute prediction module, and the downstream task head, 3) freezing all the parameters of GRIT and fine-tuning the human attribute prediction module alone. However, we found continued training as described in section 3.4 to perform best.
We used the Optuna framework Akiba et al. (2019) for hyperparameter search, closely following the experimental settings in Soni et al..
4 Experiments
We compare the performance of PLMs adapted to socio-demographic group factors, individual human contexts, and both individual and group contexts. We use the socio-demographic adapted BERT models: BERTage-MLM and BERTDS from Hung et al. (2023), HaRT Soni et al. (2022), and GRIT models, respectively. We use GPT-2HLC from Soni et al. (2022) as a PLM adapted to the social media domain but devoid of human context. We evaluate performances on two user-level regression tasks: predicting age and a personality score (openness), and on three document-level tasks: stance detection, topic detection, and age classification.
4.1 User Level Regression Tasks
We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment. These tasks require continuous outcomes derived from multiple documents written by a user. We use a subset of the data from consenting Facebook users who shared their demographic and personality scores Kosinski et al. (2012); Park et al. (2015) along with their Facebook posts. This data is essentially the same as GRIT’s pre-training data. However, the test set is from Park et al. (2015) on which HaRT and GPT-2HLC are evaluated.
GRIT is pre-trained on a multi-task learning setup, including predicting a continuous socio-demographic group attribute. GRITage is trained to predict age, and GRITope is trained to predict openness. We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively. We further evaluate these models on the test sets for personality assessment and age estimation after continuing training for these tasks, as described in section 3.4.
We use the results from Soni et al. (2022) for HaRT and GPT-2HLC which are directly comparable to GRIT models trained on the same data splits and metrics. Soni et al. fine-tuned the recurrence module of the pre-trained HaRT model for the tasks of age estimation and personality assessment using the average of user-states from non-padded blocks of texts from an author, resulting in two fine-tuned models. Similarly, they fine-tune the last two layers of the pre-trained GPT-2HLC model for these tasks. Since GPT-2HLC can not handle all text from a user in one pass, they average the predictions across all user messages corresponding to the same label for each message.
Age Estimation
The training and development datasets are identical to the pre-training data for GRIT. Age was self-reported and limited to users under 65. We compare the performance of the models on a test set of 5000 users from Park et al. (2015) and report Pearson correlation ($r$).
Personality Assessment
As for age estimation, the training and development datasets are the same as the pre-training data for GRIT. We compare the performance of the models to predict openness (one’s tendency to be open to new ideas) on a test set of 1943 users and report disattenuated Pearson correlation ($r_{dis}$) metric to account for questionnaire reliability as in Soni et al. (2022).
4.2 Document-Level Classification Tasks
We compare different models for stance detection vs. topic detection and age classification. These tasks classify a single input document (tweet message or a review) a user writes into label categories. For stance detection, we also use the historical messages of a user where available, as in Soni et al. (2022). We do not have the user information or any user historical language available for the other two tasks, so we evaluate on the single document input.
All models process the input document(s) and feed the layer-normed last non-padded token representation to the classification layer to classify the document into label categories. Only GRIT and HaRT incorporate user information and the historical language (where available). The other two models can only use the input document without a hierarchical structure to make the predictions. We compare with the results from Soni et al. (2022) and Hung et al. (2023) wherever applicable and fine-tune all the parameters of the respective pre-trained models and the classification heads for other task-model combinations using the standard cross-entropy loss.
Stance Detection
Given a single annotated tweet, this task predicts a user’s stance as in favor of, against, or neutral towards one of the five targets: atheism, climate change as a real concern, feminism, Hillary Clinton, and legalization of abortion. We fine-tune the models under comparison for each target separately. We report average of weighted F1 scores with three labels across all five targets. We use Soni et al. (2022)’s train/dev/test split over SemEval 2016 dataset (Mohammad et al., 2016). HaRT and GRIT models maintain the temporal accuracy by using only the messages posted earlier than the labeled messages from the extended dataset (Lynn et al., 2019) as a user’s historical language. We compare the results of fine-tuned GPT-2HLC, HaRT (Soni et al., 2022), and fine-tuned GRITage and GRITope.
Topic Detection
We use the US subset of the TrustPilot reviews dataset Hovy (2015) from two age groups: below 35 or above 45  [^2] As suggested by Hovy (2015), this split of the age ranges results in roughly equally-sized data sets and is non-contiguous, avoiding fuzzy boundaries.. Given a single review, this task predicts the review topics from five categories: Flights, Online marketplace, Fitness & Nutrition, Electronics, and Hotels. We use the same train, development, and test set splits as Hung et al. (2023) to eliminate any skew in the demographically-conditioned label distribution. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
Demographic Attribute Classification
We use the same subset of the TrustPilot dataset as for topic detection and the same train, development, and test splits from Hung et al. (2023). Given a single review, this task predicts the age group binary label. Age categories are equally represented in each set. We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
4.3 Human Language Modeling
To compare the effects of individual and group factors on language modeling performance, we evaluate on the test set from the pre-trained data splits. We report and compare perplexity scores from the pre-trained GPT-2 (GPT-2frozen), GPT-2HLC, HaRT, GRITage and GRITope for the human language modeling task.
5 Results and Discussion
We report results for all the tasks here, discussing their respective impacts on adapting PLMs to individual human context, group context, and both individual and group context.
5.1 Comparisons Study
User-Level Tasks.
Table 1 shows the two user-level task results. For computational reasons, both HaRT and GRIT use just 4 blocks of training data for these tasks. We find that the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality. In addition, GRIT performs better on the other user-level task using transfer learning and continued training than by fine-tuning it on the averaged user states. Comparing the transfer learning results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find that training PLMs with individual and group human context benefits such multi-document user-level regression tasks. Consequently, we may view the group attribute prediction as a regularizer for the model.
Model Human Context Age ($r$) OPE ($r_{dis}$)
GPT-2HLC
None
0.839
0.521
HaRT
Individual
0.868
0.619
GRITage
Ind + Grp
$\bf{0.890}^{\dagger}$
$\bf{0.658}^{\dagger}$
GRITope
Ind + Grp
$0.884^{\dagger}$
$0.643^{\dagger}$
[…]
Document-Level Tasks.
Model Human Context Stance (F1wtd) TD (<35) (F1mac) TD (>45) (F1mac) AC (F1mac)
GPT-2HLC
None
68.6
$\bf{69.8}$
65.4
63.9
BERTage-MLM
Group
68.4
64.6
$61.9^{*}$
BERTDS
Group
69.3
65.0
$64.1^{*}$
HaRT
Individual
$\bf{71.1}^{\dagger}$
$\bf{69.8}^{\dagger}$
65.6
$\bf{64.3}^{\dagger}$
GRITage
Ind + Grp
70.8
69.2
64.5
62.7
GRITope
Ind + Grp
70.1
66.5
64.8
61.2
[…]
Table 2 shows the results for stance detection. Both GRIT and HaRT models were fine-tuned with historical language for stance. The PLM trained with individual human context (HaRT) better detects user stance.
Table 2 also compares performances of the models on topic detection (TD) and demographic attribute classification (AC). We use GRIT models pre-trained with 8 training blocks of user texts. Both HaRT and GRIT models have a notion of the user and treat each input as written by a different user. This aspect may aid their performances even for tasks where historical language is unavailable. Even though this information is lacking for TD and AC, fine-tuned HaRT models perform better than the rest. The additional group-attribute pre-training of GRIT models may be introducing noise for document-level tasks, since the results are close to that of fine-tuned HaRT models, yet slightly lower. We can draw parallels between the performance enhancements from GPT-2HLC to HaRT, i.e., a PLM adapted to an out-of-domain corpus (social media) to a PLM trained with individual human context using the same corpus, and between the performance gains from BERTage-MLM to BERTDS, i.e., a PLM adapted to an out-of-domain corpus (blogs) to a PLM trained with group context using the same corpus. Note that Hung et al. (2023) do not report results from BERTage-MLM and BERTDS out-of-domain demographically specialized models. Our results indicate that single-document annotated classification tasks may benefit simply by training PLMs with individual human context.
Perplexity.
We also compare the language modeling capability of the various models. Table 3 reports perplexity on the test set of 4438 users from the pre-training data. Multi-task learning is known to impact individual task performance, so we expect a slight dip in perplexity. The results align with our hypothesis and the trend shown in Soni et al. (2022). The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT models perform best while GRIT models result in a slightly lower perplexity than HaRT.
Model Human Context
Test ($ppl$)
GPT-2frozen
None
114.82
GPT-2HLC
None
36.39
HaRT
Individual
28.24
GRITage
Ind + Grp
31.77
GRITope
Ind + Grp
30.32
Table 3: Comparing perplexity on language modeling for models trained with individual and group contexts.
5.2 Error Analysis and Disparity
We perform a set of error analysis by comparing performance metrics of HaRT and GRIT models (pre-trained with 4 training blocks) for the user-level regression tasks of age and openness prediction across different groups based on a demographic factor. The different groups are created by sampling the test set into the following age buckets: below 18, 18-21, 21-30, 30-45, and above 45. Table 4 shows GRITage performs better for the task of estimating age across all age groups i.e., exhibits lesser error. We also see lesser errors in GRIT models for the openness assessment tasks (Appendix Table 6) as well as conforming results on both tasks when comparing using the MSE metric (Appendix Tables 7 and 8).
Additionally, we use the error analysis results to compare the error disparity Shah et al. (2020) in GRIT models and HaRT. Error disparity can be exemplified by the "Wall Street Journal Effect" – a systematic difference in error as a function of demographics Hovy and Søgaard (2015). It can be calculated as the difference in the computed metric across different groups based on a demographic factor Shah et al. (2020). We compute the mean error disparity ($MED$) as the sum of the differences in the metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) computed for each group averaged by the number of difference pairs.
Table 8 reports the $MED$ for each model-task pair for HaRT, GRITage, and GRITope models, and age estimation and openness assessment tasks. We find GRIT models to demonstrate lower mean error disparity for each metric i.e., making less error as a function of the age groups.
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
503
0.223
0.394
0.393
18-21
560
0.230
0.278
0.276
21-30
563
0.512
0.531
0.519
30-45
249
0.485
0.530
0.520
45+
68
0.106
0.205
0.180
Table 4: Pearson $r$ for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row.
[…]
Age ($r$)
0.215
0.181
0.185
OPE ($r_{dis}$)
0.075
0.090
0.072
Table 5: Mean error disparity for age over five age buckets. Bold indicates best in column (lower is better).
6 Related Work
People use language to communicate and convey meaning more than mere words. Much work in human-centered NLP has focused on identifying and evaluting including human context in our models. Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups. These works improved sentiment analysis, polarity classification, and topic detection. Other researchers explored human factors like social networks (Huang et al., 2014; Radfar et al., 2020), occupation (Preoţiuc-Pietro et al., 2015), personality (Schwartz et al., 2013; Lynn et al., 2017), and social media attributes (Bamman and Smith, 2015; Lynn et al., 2019) to improve toxic language detection, sarcasm detection, or stance detection.
Some studies go beyond explicit groups and learning individual representations latently or via historical language. Jaech and Ostendorf (2018) learned latent user embeddings for search query completion. Delasalles et al. (2019) conditioned a language model on a recurrently updated latent author representation. Welch et al. (2020) motivate personalized word embeddings by jointly learning a latent representation for each user and generic word representations. Hofmann et al. (2021) combined learned latent representations of the social space with time to produce dynamic contextualized word embeddings. King and Cook (2020) created personalized models using the authors’ historical text. Lynn et al. (2020) attend to user’s past messages to better predict user personality.
Research on adapting pre-trained language models to socio-demographic factors has been minimal. Guda et al. (2021) propose EMPATH-BERT, a demographically-aware model to predict empathy and distress better. Lauscher et al. (2022) probe PLMs to understand if their representations encode socio-demographic information. Hung et al. (2023) generalize the task-specific EMPATH-BERT to create a PLM injected with demographic group information using a dynamic multi-task learning setup. We adapt their mono-lingual BERT-based model to age with out-of-domain data for our comparison.
Several studies (Li et al., 2021; Mireshghallah et al., 2022; Zhong et al., 2021) have explored adapting pre-trained Transformer-based language models to individual human contexts for downstream tasks. Li et al. (2021) find benefits in adding user IDs to the input to generate explanations for recommender systems. Zhong et al. (2021) learned a latent user-specific vector prepended to the input embeddings to classify sentiment better, similar to Mireshghallah et al. (2022), who use static user text identifiers instead. Matero et al. (2021) perform masked language modeling on users’ past messages with message-level attention, producing efficient document representations for stance detection. Soni et al. (2022) propose human language modeling, where language is modeled conditioned on a dynamic user state derived from temporally ordered past user utterances. We use their state-of-the-art model HaRT in our comparison and as a base for our pre-trained model with individual and group human context.
7 Conclusion
NLP benefits from modeling group traits like sociodemographic factors and individual users in terms of latent human context. However, humans exhibit varying degrees of group and individual characteristics. Through a comparative study of five user- and document-level tasks, we uncover how using individual traits and group characteristics in PLMs optimizes user-level regression tasks like age estimation and openness assessment. Meanwhile, individual human context training alone appears to bolster single-document annotated classification tasks like stance and topic detection. Despite our progress, our research reveals there are still considerable strides to be made in modeling human factors in language models. Our findings provide valuable insight into including human context in pre-trained language models to suit specific applications.
Limitations
The purpose of our study is to compare the impacts of modeling sociodemographic group attributes and modeling individual user traits, and we use relevant models to represent each of the approaches. There are likely to be other ways to model these approaches and the models we use are only one of the ways. Additionally, these models in themselves have limitations like the blocks mechanism to process all the text from author induces compute requirements resulting in a capping of the number of blocks used for training. While it is also unclear how many blocks are sufficient to capture the human context, and if it is helpful to use the earliest language or the most recently used language in the capped number of blocks.
Secondly, some of the datasets (TrustPilot) used do not have appropriate user identification or historical language to create an individual human context.
Lastly, as noted earlier, models and data that touch upon sensitive user information require an extremely responsible usage and limit researchers to make them publicly available.
Ethical Considerations
Models that incorporate sociodemographic information need to be considered with special scrutiny. On the one hand, they have the potential to produce fairer and more inclusive results, because they can account for human language variation. On the other hand, they risk revealing identifying or sensitive information, which can lead to profiling and stereotyping. These may present opportunities for unintended malicious exploitations. For example, models that improve demographic groups prediction or psychological assessments could be used for targeting content for individuals without their awareness or consent. Such models may also risk release of private information of the research participant if trained on private data unchecked for exposing identifying information. For this reason, we take a conservative release strategy. While we support open research and reproducibility, data and privacy protection take precedence. Thus, we will only be releasing the code for our comparison study and the data that does not contain sensitive information i.e., stance detection datasets and TrustPilot datasets for topic detection and attribute classification. This is also in accordance with the DUA we have received from the authors of the papers/models that we employ in our work.
Our comparison study aims to guide and further speed the growing body of human-centered AI research. The models under comparison aim to enable applicability in the interdisciplinary studies of the human condition leading to helpful tools for psychological health. However, at this point these models are not intended for use in practice and should be evaluated for failures. All user-level tasks presented here were reviewed and approved or exempted by an academic institutional review board (IRB). Our studies are limited to US-English due to comparability reasons. However, similar effects are likely to hold for other languages, and should be evaluated in future work.
Appendix A Appendix
A.1 Experimental Settings
We closely follow the experimental settings from Soni et al. (2022) and similarly use Optuna framework (Akiba et al., 2019) for hyperparameter search. We search for learning rates between 5e-6 and 5e-4, and between 1e-7 and 1e-5 for different tasks. We will make our best found hyperparameter values publicly available with our code and results in the github repository. All experiments are run on NVIDIA RTX A6000 GPUs of 48GB. Pre-training takes approx 14 hours for 1 epoch and fine-tuning takes approx 1-4 hours depending on the task.
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
503
0.627
0.644
0.618
18-21
560
0.557
0.608
0.592
21-30
563
0.715
0.741
0.738
30-45
249
0.594
0.669
0.667
45+
68
0.567
0.546
0.599
Table 6: Disattenuated pearson $r$ for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row.
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
1113
4.07
2.52
2.82
18-21
1387
6.52
4.00
3.89
21-30
1557
17.82
12.64
13.11
30-45
695
48.59
39.79
40.43
45+
248
114.92
121.66
134.72
Table 7: Mean squared error for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).
Age bucket
#Users
HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
<18
503
0.423
0.410
0.429
18-21
560
0.496
0.487
0.506
21-30
563
0.429
0.380
0.381
30-45
249
0.578
0.489
0.489
45+
68
0.584
0.501
0.467
Table 8: Mean squared error for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).

### v2

@@
-Comparing Human-Centered Language Modeling: Is it Better to Model Groups, Individual Traits, or Both?
+Comparing Pre-trained Human Language Models: Is it Better with Human Context as Groups, Individual Traits, or Both?
 Abstract
-Natural language processing has made progress in incorporating human context into its models, but whether it is more effective to use group-wise attributes (e.g., over-45-year-olds) or model individuals remains open.
-Group attributes are technically easier but coarse: not all 45-year-olds write the same way.
-In contrast, modeling individuals captures the complexity of each person’s identity.
-It allows for a more personalized representation, but we may have to model an infinite number of users and require data that may be impossible to get.
-We compare modeling human context via group attributes, individual users, and combined approaches.
-Combining group and individual features significantly benefits user-level regression tasks like age estimation or personality assessment from a user’s documents.
-Modeling individual users significantly improves the performance of single document-level classification tasks like stance and topic detection.
-We also find that individual-user modeling does well even without user’s historical data.
-Comparing Human-Centered Language Modeling:
-Is it Better to Model Groups, Individual Traits, or Both?
-[…]
+Incorporating human context into language models is the next frontier for human-centered natural language processing.
+Currently, two pre-training methods exist: group-wise attributes (e.g., over-45-year-olds) or individual traits.
+Group attributes are coarse — not all 45-year-olds write the same way — while modeling individual traits allows for a more personalized representation, but requires more complex modeling and data.
+So far, it is unclear which pre-training approach benefits what tasks.
+We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on 5 user- and document-level tasks.
+We find that pre-training with both group and individual features significantly improves the two user-level regression tasks like age estimation and personality assessment.
+Pre-training on individual users significantly improves the three document-level classification tasks like stance and topic detection.
+It even does well for downstream tasks without historical user data.
+Our results suggest both approaches have specific use cases, opening new avenues for human-centered language modeling.
+Comparing Pre-trained Human Language Models:
+Is it Better with Human Context as Groups, Individual Traits, or Both?
+Nikita Soni1, Niranjan Balasubramanian1, H. Andrew Schwartz1,  and Dirk Hovy2 1Stony Brook University, 2Bocconi University {nisoni, niranjan, has}@cs.stonybrook.edu, dirk.hovy@unibocconi.it
 1 Introduction
 Language varies between people.
-Two strands of human-centered NLP work have modeled the humans behind the language.
-The first focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group.
-These attributes include socio-demographics like age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more.
-The second strand focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
-While these two strands have advanced human-centered NLP (Soni et al., 2023), their relative strengths, complementarity, and impact over different tasks are poorly understood.
-People are not defined by their group membership alone Orlikowski et al.
+To capture this notion, two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language.
+The first approach focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group.
+These socio-demographic attributes include age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more.
+The second approach focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
+Figure 1: Pre-training a language model with no human context, with socio-demographic group attribute, with individual traits, and with both group and individual traits.
+While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2023): People are not defined by their group membership alone Orlikowski et al.
 (2023), but individual traits might not generalize.
-This paper compares the two approaches and their combination in the same framework and tasks in pre-trained large language models.
+We compare the downstream performance of models pre-trained with the two approaches as well as their combination.
+We answer the following broader research questions:
+(RQ1): How can we incorporate group and/or individual human context into pre-training?
+(RQ2): How does pre-training with different human contexts differ in terms of downstream performance for different tasks?
 We use the monolingual socio-demographically adapted model from Hung et al.
 (2023) and the HaRT model from Soni et al.
-(2022) for the first two types.
-We use a multi-task learning setup to create GRIT, a PLM based on HaRT, trained with both individual and group human context in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness).
+(2022) for group and individual human contexts, respectively.
+For a model trained with both individual and group human context, we use a multi-task learning PLM based on HaRT to create a model, GRIT, in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness).
 We test all systems on five user- and document-level tasks.
-Note that because we focus on conceptually comparing group and individual traits, we cannot compare to GPT4, which does not support stratification to either attribute.
-PLMs trained on individuals and groups enhance user-level regression tasks like age estimation and personality assessment from user’s multiple documents.
-Such user-level tasks focus on individual person, and our findings show that these people are best modeled as a combination of their groups and individual traits conforming with the notion that a person is a mix of their group attributes and unique characteristics.
-Document-level categorization tasks, like stance detection, are more personal.
-A PLM taught within an individual human context alone improves our considered document-level classification tasks.
+Note that because we focus on empirically comparing pre-training group and individual traits in PLMs, we cannot compare to large language models like GPT4, which do not support stratification to either attribute via pre-training.
+Recent studies have explored methodologies like user-adapters (Zhong et al., 2021), and user-centric prompting (Li et al., 2023) to include human context into models.
+In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts.
+We find that PLM pre-trained on individuals and group features enhances 2 user-level regression tasks: age estimation and personality assessment from multiple user documents.
+These user-level tasks focus on individual people.
+Our findings, that they are best modeled as a combination of their group attributes and unique characteristics, conforms with the notion that a person is a mix of both.
+Document-level categorization tasks, like stance detection, are more personal than group based.
+A PLM pre-trained on individual human context alone improves 3 document-level classification tasks.
 By their very nature, models of this kind touch upon sensitive user information.
@@
 For more information about the model and data, see Sections 3 and 4.
-For a discussion of the ethical implications of the model and data, see Section Ethical Considerations.
+For a discussion of the ethical implications of the model and data, see Ethical Considerations section.
 Contributions.
-Our contributions are: (1) We provide an analysis with a comparison of modeling individual users, group socio-demographic features, and both group and individual traits in PLMs.
-(2) We evaluate the three modeling strategoies on five downstream tasks: two user-level (personality-openness evaluation and age estimation) and three document-level classification tasks (stance detection, topic detection, and age category prediction).
-(3) We find that user-level regression tasks like estimating age and assessing personality from user’s multiple documents perform better with mixed individual and group human context (GRIT) and document-level tasks like stance detection and topic detection perform better with individual context alone (HaRT).
+(1) We compare pre-training several language models with individual users, group socio-demographic features, and both traits.
+(2) We evaluate the three specific pre-training strategies on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category prediction).
+(3) We find that the two user-level regression tasks perform better when pre-trained with both strategies (GRIT) and the three single document-level tasks perform better on fine-tuning a language model pre-trained with individual context alone (HaRT).
 2 Integrating Human Context in PLMs
-For our comparative study, we use three systems representing the paradigms of human context modeling to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts.
-Training with group context.
+For our comparison, we use three systems representing the paradigms of pre-training with human context (Figure 1).
+We want to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts.
+As noted earlier, we focus on recent approaches for pre-training language models with the additional human context.
+Pre-training with group context.
 We build on Hung et al.
-(2023)’s work to explore demographic adaptation in transformer-based PLMs.
+(2023), a recent model to explore demographic adaptation in transformer-based PLMs.
 They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author.
@@
 They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection.
-In our study, we use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4).
-We use the results from Hung et al.
-(2023) with the monolingual PLM model BERT on a US-based English dataset with out-of-domain demographic (age) specialization for our comparison study.
+We use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4).
+We compare to the results from Hung et al.
+(2023) with monolingual BERT on a US-English dataset with out-of-domain demographic (age) specialization.
 Out-of-domain data is the Blogs authorship corpus Schler et al.
 (2006), and in-domain means Trustpilot corpus.
-To be consistent and fair in comparing with other human context training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
-Training with individual human context.
+To be consistent and fair in comparing with other human context pre-training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
+Pre-training with individual human context.
 Soni et al.
-(2022) introduced human language modeling (HuLM), i.e., training a regular language modeling task but including a dynamic individual human context vector derived from the authors’ texts.
-This vector captures the human states in which the text is generated to induce coherence between different texts generated by the same author.
-Soni et al.
-also view this vector as representing the text-derived human factors.
-They introduce a Human-aware Recurrent Transformer (HaRT), a unidirectional autoregressive PLM that trains for the HuLM task.
+(2022) introduced human language modeling (HuLM) in pre-trained LMs.
+I.e., pre-training a regular language modeling but including a dynamic individual human context vector derived from the authors’ historical texts.
+This vector captures the human states in which the text was generated.
+It also adds coherence to texts generated by the same author.
+They introduce a Human-aware Recurrent Transformer (HaRT), an autoregressive PLM for human language modeling.
 They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks.
 We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
-Training with both group and individual human context.
-We train a PLM that can integrate the author’s individual and group human context knowledge when training for language modeling.
-We extend Soni et al.’s HaRT model by training for HuLM and predicting a human group attribute in a dynamic muti-task learning setup as used by Hung et al..
-We want to induce the individual human context through the author’s language and inject the group context by predicting a group attribute of the author.
-Predicting group attributes during training can also be viewed as a regularizer for the model, as it constrains the possible output.
-We discuss the model in detail in section 3.
-We compare two user-level tasks and a document-level task, as in Soni et al.
-(2022), and on topic detection and age prediction tasks, as in Hung et al.
-(2023).
+Pre-training with both group and individual human context.
+We train a PLM to integrate both individual and group human context in language modeling pre-training.
+We extend HaRT by predicting an additional human group attribute in a dynamic muti-task learning setup similar to Hung et al..
+To facilitate comparison, we combine the pre-training inducing the individual human context through the author’s language and inject the group context by predicting a group attribute of the author.
+We choose downstream tasks from both prior works (Soni et al., 2022; Hung et al., 2023) to compare our three language models pre-trained with different levels of human contexts: two user-level regression tasks and three single document-level classification task.
 3 Models
-This section describes the models we compare to represent group and individual human contexts.
-3.1 Modeling individual human context
+3.1 Pre-training with individual human context
 HaRT.
@@
 $\prod_{a\in\textup{Users}}\prod_{t=1}^{|\mathcal{B}_{a}|}\prod_{i=1}^{|B_{t}^{(a)}|}Pr(w_{t,i}|w_{t,1:i-1},B_{1:t-1}^{(a)})$
-where, $w_{t,i}$ is the $i^{th}$ token in the $t^{th}$ block ($B_{t}^{(a)}$) for user $a$. The tokens from the previous blocks are represented using HaRT’s recurrently updated user-state vector.
-3.2 Modeling group human context
+where, $w_{t,i}$ is the $i^{th}$ token in the $t^{th}$ block ($B_{t}^{(a)}$) for user $a$.
+The tokens from the previous blocks are represented using HaRT’s recurrently updated user-state vector.
+Soni et al.
+use cross-entropy loss for the HuLM objective.
+3.2 Pre-training with group human context
 BERTDS and BERTage-MLM.
@@
 (2018) of losses, they adopt a dynamic MTL objective for training with group human context.
-Kendall et al.
-interpret homoscedastic uncertainty as task-dependent weighting and derive a multi-task loss function that can optimally learn the weights to balance the impact of multiple loss functions.
+Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions.
 This approach accounts for the different scales across multiple losses.
@@
 Hung et al. minimize the sum of both the uncertainty adjusted losses: $\tilde{L_{mlm}}+\tilde{L_{dem}}$.
-3.3 Modeling both individual and group human context
+3.3 Pre-training with both individual and group human context
 GRIT.
-GRIT incorporates both individual and group human contexts using a multi-task learning setup with HaRT that also predicts a socio-demographic attribute of the author in each forward pass using the average of user-state vectors from each non-padded block of the user’s temporally ordered text.
-Multi-Task Learning.
-GRIT is pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup.
+GRoup and Individual HaRT is a LM pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup.
 The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
@@
 This representation is layer-normed and linearly transformed before making a continuous-valued prediction for the specific attribute.
-We train one model for the attribute age (GRITage) and one for the attribute personality type openness (GRITope), respectively.
+We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope), respectively.
 The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$).
-To account for the homoscedastic uncertainty Kendall et al.
-(2018) of losses, we adopt a dynamic MTL objective as Hung et al.
-(2023).
+Similar to Hung et al.
+(2023), we account for the homoscedastic uncertainty Kendall et al.
+(2018) of losses.
 We use the joint loss for a continuous and discrete output as derived in Kendall et al.
@@
 Since HaRT caps to 4 train blocks for user-level downstream tasks, we also pre-train GRITage and GRITope with four training blocks.
-3.4 Transfer Learning
-To assess the efficiency of GRIT to transfer learning from predicting one group human attribute to another, we experiment with continuing pre-training for each group attribute.
-To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness).
-We do the same for GRITope and continued MTL with predicting age.
 3.4 Fine-Tuning
@@
 We used the Optuna framework Akiba et al. (2019) for hyperparameter search, closely following the experimental settings in Soni et al..
+3.5 Transfer Learning
+We experiment with continuing pre-training GRIT models for each group attribute.
+To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness).
+We do the same for GRITope and continued MTL with predicting age.
 4 Experiments
-We compare the performance of PLMs adapted to socio-demographic group factors, individual human contexts, and both individual and group contexts.
-We use the socio-demographic adapted BERT models: BERTage-MLM and BERTDS from Hung et al.
-(2023), HaRT Soni et al.
-(2022), and GRIT models, respectively.
-We use GPT-2HLC from Soni et al.
+Our research goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined.
+To this end, we evaluate performances of the models defined in Section 3 on two user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification.
+We also compare against GPT-2HLC from Soni et al.
 (2022) as a PLM adapted to the social media domain but devoid of human context.
-We evaluate performances on two user-level regression tasks: predicting age and a personality score (openness), and on three document-level tasks: stance detection, topic detection, and age classification.
 4.1 User Level Regression Tasks
-We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment.
-These tasks require continuous outcomes derived from multiple documents written by a user.
-We use a subset of the data from consenting Facebook users who shared their demographic and personality scores Kosinski et al.
+We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment Kosinski et al.
 (2013); Park et al.
-(2015) along with their Facebook posts.
-This data is essentially the same as GRIT’s pre-training data.
-However, the test set is from Park et al.
-(2015) on which HaRT and GPT-2HLC are evaluated.
-GRIT is pre-trained on a multi-task learning setup, including predicting a continuous socio-demographic group attribute.
-GRITage is trained to predict age, and GRITope is trained to predict openness.
+(2015).
+These social scientific tasks require predicting continuous outcomes for a user given multiple documents written by them.
+We use the same data splits as used by Soni et al.
+(2022) for evaluating HaRT and GPT-2HLC
 We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively.
-We further evaluate these models on the test sets for personality assessment and age estimation after continuing training for these tasks, as described in section 3.4.
+We further evaluate these models on the test sets for personality assessment and age estimation, respectively, after continuing training for these tasks, as described in section 3.5.
 We use the results from Soni et al.
@@
 Since GPT-2HLC can not handle all text from a user in one pass, they average the predictions across all user messages corresponding to the same label for each message.
-Age Estimation
-The training and development datasets are identical to the pre-training data for GRIT.
-Age was self-reported and limited to users under 65.
-We compare the performance of the models on a test set of 5000 users from Park et al.
-(2015) and report Pearson correlation ($r$).
-Personality Assessment
-As for age estimation, the training and development datasets are the same as the pre-training data for GRIT.
-We compare the performance of the models to predict openness (one’s tendency to be open to new ideas) on a test set of 1943 users and report disattenuated Pearson correlation ($r_{dis}$) metric to account for questionnaire reliability as in Soni et al.
-(2022).
 4.2 Document-Level Classification Tasks
@@
 (2023).
-Given a single review, this task predicts the age group binary label.
+Given a single review, this task predicts the age group binary label (<35 years old or >45 years old).
 Age categories are equally represented in each set.
@@
 5 Results and Discussion
-We report results for all the tasks here, discussing their respective impacts on adapting PLMs to individual human context, group context, and both individual and group context.
+We report results for all the tasks here, discussing their respective impacts from pre-training LMs with individual human context, group context, and both individual and group context.
 5.1 Comparisons Study
-User-Level Tasks.
-Table 1 shows the two user-level task results.
-For computational reasons, both HaRT and GRIT use just 4 blocks of training data for these tasks.
-We find that the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality.
-In addition, GRIT performs better on the other user-level task using transfer learning and continued training than by fine-tuning it on the averaged user states.
-Comparing the transfer learning results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find that training PLMs with individual and group human context benefits such multi-document user-level regression tasks.
-Consequently, we may view the group attribute prediction as a regularizer for the model.
+User-Level Regression Tasks.
+Table 1 shows the two user-level regression task results.
+We find that just the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality.
+Additionally, comparing the transfer learning (Section 3.5) results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find the GRIT models to fare better.
+Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it is still devoid of human context.
+HaRT adds pre-training with the individual human context, and GRIT adds pre-training with both group and individual human contexts (Figure 1).
+As Table 1 shows, there are gains going from GPT-2HLC (no human context) to HaRT (individual human context) and further to GRIT (individual + group human context).
+This indicates that pre-training PLMs with individual and group human context may benefit multi-document user-level regression tasks like the ones we considered.
 Model Human Context Age ($r$) OPE ($r_{dis}$)
-GPT-2HLC
-None
-0.839
-0.521
-HaRT
-Individual
-0.868
-0.619
-GRITage
-Ind + Grp
-$\bf{0.890}^{\dagger}$
-$\bf{0.658}^{\dagger}$
-GRITope
-Ind + Grp
-$0.884^{\dagger}$
-$0.643^{\dagger}$
+GPT-2HLC None 0.839 0.521
+HaRT Individual 0.868 0.619
+GRITage Ind + Grp 0.890 0.658
+GRITope Ind + Grp $0.884$ $0.643$
+Table 1: Pearson $r$ for age, disattenuated Pearson $r$ for openness.
+Pre-training with individual plus group context show benefits in estimating age and assessing personality (openness).
+Bold = best in column.
+We find no statistical difference between GRITage and GRITope for the task of age estimation.
+All other results show statistical significance $p<0.05$ using paired t-test.
+Document-Level Classification Tasks.
+Model Human Context Stance (F1wtd) TD (<35) (F1mac) TD (>45) (F1mac) AC (F1mac)
+GPT-2HLC None 68.6 69.8 65.4 63.9
+BERTage-MLM Group - 68.4 64.6 $61.9$
+BERTDS Group - 69.3 65.0 $64.1$
+HaRT Individual $\textbf{71.1}^{*}$ $\textbf{69.8}^{*}$ 65.6 $\textbf{64.3}^{*}$
+GRITage Ind + Grp 70.8 69.2 64.5 62.7
+GRITope Ind + Grp 70.1 66.5 64.8 61.2
 […]
-Document-Level Tasks.
-Model Human Context Stance (F1wtd) TD (<35) (F1mac) TD (>45) (F1mac) AC (F1mac)
-GPT-2HLC
-None
-68.6
-$\bf{69.8}$
-65.4
-63.9
-BERTage-MLM
-Group
-68.4
-64.6
-$61.9^{*}$
-BERTDS
-Group
-69.3
-65.0
-$64.1^{*}$
-HaRT
-Individual
-$\bf{71.1}^{\dagger}$
-$\bf{69.8}^{\dagger}$
-65.6
-$\bf{64.3}^{\dagger}$
-GRITage
-Ind + Grp
-70.8
-69.2
-64.5
-62.7
-GRITope
-Ind + Grp
-70.1
-66.5
-64.8
-61.2
-[…]
-Table 2 shows the results for stance detection.
-Both GRIT and HaRT models were fine-tuned with historical language for stance.
-The PLM trained with individual human context (HaRT) better detects user stance.
-Table 2 also compares performances of the models on topic detection (TD) and demographic attribute classification (AC).
-We use GRIT models pre-trained with 8 training blocks of user texts.
-Both HaRT and GRIT models have a notion of the user and treat each input as written by a different user.
-This aspect may aid their performances even for tasks where historical language is unavailable.
-Even though this information is lacking for TD and AC, fine-tuned HaRT models perform better than the rest.
-The additional group-attribute pre-training of GRIT models may be introducing noise for document-level tasks, since the results are close to that of fine-tuned HaRT models, yet slightly lower.
-We can draw parallels between the performance enhancements from GPT-2HLC to HaRT, i.e., a PLM adapted to an out-of-domain corpus (social media) to a PLM trained with individual human context using the same corpus, and between the performance gains from BERTage-MLM to BERTDS, i.e., a PLM adapted to an out-of-domain corpus (blogs) to a PLM trained with group context using the same corpus.
-Note that Hung et al.
-(2023) do not report results from BERTage-MLM and BERTDS out-of-domain demographically specialized models.
-Our results indicate that single-document annotated classification tasks may benefit simply by training PLMs with individual human context.
+Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute classification (AC).
+We see that task fine-tuned HaRT (individual human context) models performs better on all tasks..
+HaRT models inherently have an additional context of the individual user and do not treat all inputs as if written by the same user.
+The tasks considered here relate more to personal opinions and preferences, rather than group-level ones.
+HaRT model is well-suited for incorporating such personalization due to its pre-training with individual human context.
+Additionally, the results indicate that the models pre-trained with group context (BERTDS) fare well in the group-based tasks of topic detection and age classification.
+Whereas, the models pre-trained on both individual and group human context (GRIT) appears to bring in noise for the group-based and personal stance detection tasks resulting in a slightly worse performance.
+Further, it is important to note that the individual human context (HaRT) derived from the historical tweets from the users in the stance detection dataset provides for a richer human context.
+Although, the performance is not greatly hurt even if historical language is not available for certain tasks (TD and AC).
 Perplexity.
 We also compare the language modeling capability of the various models.
-Table 3 reports perplexity on the test set of 4438 users from the pre-training data.
-Multi-task learning is known to impact individual task performance, so we expect a slight dip in perplexity.
-The results align with our hypothesis and the trend shown in Soni et al.
+Table 3 reports perplexity on the test set of 4438 users.
+The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT (individual human context) models perform best while GRIT (individual + group human context) models result in a slightly lower perplexity than HaRT.
+GRIT models are pre-trained under a multi-task learning setup which most likely is hurting the individual task performances, thus resulting in a slight dip in the perplexity compared to HaRT.
+Further, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al.
 (2022).
-The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT models perform best while GRIT models result in a slightly lower perplexity than HaRT.
-Model Human Context
-Test ($ppl$)
-GPT-2frozen
-None
-114.82
-GPT-2HLC
-None
-36.39
-HaRT
-Individual
-28.24
-GRITage
-Ind + Grp
-31.77
-GRITope
-Ind + Grp
-30.32
+Model Human Context Test ($ppl$)
+GPT-2frozen None 114.82
+GPT-2HLC None 36.39
+HaRT Individual 28.24
+GRITage Ind + Grp 31.77
+GRITope Ind + Grp 30.32
 Table 3: Comparing perplexity on language modeling for models trained with individual and group contexts.
 5.2 Error Analysis and Disparity
-We perform a set of error analysis by comparing performance metrics of HaRT and GRIT models (pre-trained with 4 training blocks) for the user-level regression tasks of age and openness prediction across different groups based on a demographic factor.
-The different groups are created by sampling the test set into the following age buckets: below 18, 18-21, 21-30, 30-45, and above 45.
-Table 4 shows GRITage performs better for the task of estimating age across all age groups i.e., exhibits lesser error.
-We also see lesser errors in GRIT models for the openness assessment tasks (Appendix Table 6) as well as conforming results on both tasks when comparing using the MSE metric (Appendix Tables 7 and 8).
-Additionally, we use the error analysis results to compare the error disparity Shah et al.
-(2020) in GRIT models and HaRT.
-Error disparity can be exemplified by the "Wall Street Journal Effect" – a systematic difference in error as a function of demographics Hovy and Søgaard (2015).
-It can be calculated as the difference in the computed metric across different groups based on a demographic factor Shah et al.
-(2020).
-We compute the mean error disparity ($MED$) as the sum of the differences in the metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) computed for each group averaged by the number of difference pairs.
-Table 8 reports the $MED$ for each model-task pair for HaRT, GRITage, and GRITope models, and age estimation and openness assessment tasks.
-We find GRIT models to demonstrate lower mean error disparity for each metric i.e., making less error as a function of the age groups.
-Age bucket
-#Users
-HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
-<18
-503
-0.223
-0.394
-0.393
-18-21
-560
-0.230
-0.278
-0.276
-21-30
-563
-0.512
-0.531
-0.519
-30-45
-249
-0.485
-0.530
-0.520
-45+
-68
-0.106
-0.205
-0.180
-Table 4: Pearson $r$ for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row.
+We conduct an error analysis as a function of a socio-demographic group attribute.
+Specifically, we measure GRIT and HaRT for error disparity (Shah et al., 2020) – – a systematic difference in error as a function of demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015).
+We analyzed outcome and error disparity on age and openness prediction from both the models with individual context (HaRT) and combined individual plus group context (GRIT).
+First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set.
+Then, we compare the performance of our models across these buckets.
+Results from table 4 indicate that pre-training with individual and group context together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age.
+We see similar trends for assessing openness personality (Appendix Tables 6, 8).
+This suggests that the group attribute prediction maybe acting as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
+To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, that is averaged by the number of pairs (Shah et al., 2020).
+A lower averaged sum of differences implies lesser errors as a function of the age groups.
+Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
+Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
+<18 1113 0.223 0.394 0.393
+18-21 1387 0.230 0.278 0.276
+21-30 1557 0.512 0.531 0.519
+30-45 695 0.485 0.530 0.520
+45+ 248 0.106 0.205 0.180
+Table 4: Pearson $r$ for age over five age buckets using different types of human contexts for error analysis.
+Bold indicates best in row.
+We find no statistical difference between GRITage and GRITope for buckets 21-30 and 30-45.
+All other results show statistical significance $p<0.05$ using paired t-test.
 […]
-Age ($r$)
-0.215
-0.181
-0.185
-OPE ($r_{dis}$)
-0.075
-0.090
-0.072
-Table 5: Mean error disparity for age over five age buckets. Bold indicates best in column (lower is better).
+Age ($r$) 0.215 0.181 0.185
+OPE ($r_{dis}$) 0.075 0.090 0.072
+Table 5: Mean error disparity for age estimation and openness personality assessment over five age buckets. Bold indicates best in column (lower is better).
 6 Related Work
-People use language to communicate and convey meaning more than mere words.
-Much work in human-centered NLP has focused on identifying and evaluting including human context in our models.
+Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models.
 Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups.
 These works improved sentiment analysis, polarity classification, and topic detection.
-Other researchers explored human factors like social networks (Huang et al., 2014; Radfar et al., 2020), occupation (Preoţiuc-Pietro et al., 2015), personality (Schwartz et al., 2013; Lynn et al., 2017), and social media attributes (Bamman and Smith, 2015; Lynn et al., 2019) to improve toxic language detection, sarcasm detection, or stance detection.
-Some studies go beyond explicit groups and learning individual representations latently or via historical language.
-Jaech and Ostendorf (2018) learned latent user embeddings for search query completion.
-Delasalles et al.
-(2019) conditioned a language model on a recurrently updated latent author representation.
-Welch et al.
-(2020) motivate personalized word embeddings by jointly learning a latent representation for each user and generic word representations.
-Hofmann et al.
-(2021) combined learned latent representations of the social space with time to produce dynamic contextualized word embeddings.
-King and Cook (2020) created personalized models using the authors’ historical text.
-Lynn et al.
-(2020) attend to user’s past messages to better predict user personality.
-Research on adapting pre-trained language models to socio-demographic factors has been minimal.
-Guda et al.
-(2021) propose EMPATH-BERT, a demographically-aware model to predict empathy and distress better.
-Lauscher et al.
-(2022) probe PLMs to understand if their representations encode socio-demographic information.
+Other researchers explored adapting to user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) to improve downstream tasks like sarcasm detection, and toxic language detection.
+Some studies go beyond explicit groups and learning individual representations latently Jaech and Ostendorf (2018); Delasalles et al.
+(2019) or via historical language Matero et al.
+(2021).
+With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks.
+More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context.
+However, few studies have explored including human context within the pre-training regime of LMs.
 Hung et al.
-(2023) generalize the task-specific EMPATH-BERT to create a PLM injected with demographic group information using a dynamic multi-task learning setup.
-We adapt their mono-lingual BERT-based model to age with out-of-domain data for our comparison.
-Several studies (Li et al., 2021; Mireshghallah et al., 2022; Zhong et al., 2021) have explored adapting pre-trained Transformer-based language models to individual human contexts for downstream tasks.
-Li et al.
-(2021) find benefits in adding user IDs to the input to generate explanations for recommender systems.
-Zhong et al.
-(2021) learned a latent user-specific vector prepended to the input embeddings to classify sentiment better, similar to Mireshghallah et al.
-(2022), who use static user text identifiers instead.
-Matero et al.
-(2021) perform masked language modeling on users’ past messages with message-level attention, producing efficient document representations for stance detection.
-Soni et al.
-(2022) propose human language modeling, where language is modeled conditioned on a dynamic user state derived from temporally ordered past user utterances.
-We use their state-of-the-art model HaRT in our comparison and as a base for our pre-trained model with individual and group human context.
+(2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup.
+Further, Soni et al.
+(2022) pre-train a LM with individual human context derived from user language.
+Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
 7 Conclusion
-NLP benefits from modeling group traits like sociodemographic factors and individual users in terms of latent human context.
+NLP benefits from modeling latent human context, such as sociodemographic group traits or individual tendencies.
+A recent development has been to incorporate this additional human context into the PLMs’ pre-training regimen.
 However, humans exhibit varying degrees of group and individual characteristics.
-Through a comparative study of five user- and document-level tasks, we uncover how using individual traits and group characteristics in PLMs optimizes user-level regression tasks like age estimation and openness assessment.
-Meanwhile, individual human context training alone appears to bolster single-document annotated classification tasks like stance and topic detection.
-Despite our progress, our research reveals there are still considerable strides to be made in modeling human factors in language models.
-Our findings provide valuable insight into including human context in pre-trained language models to suit specific applications.
+Learning about the impacts of pre-training with different types of human context will propel us forward in including human context in our base LLMs (Soni et al., 2023).
+To assess its impacts, we compare three types of PLMs pre-trained with sociodemographic group attributes, individual human contexts, or a combination across five user- and document-level tasks.
+We find that pre-training with individual and group human context improves the two user-level tasks: age and personality prediction.
+Pre-training with individual human context improves the three single-document classification tasks, including stance and topic detection.
+Pre-training solely on group context helps group-based document classification tasks, though not optimally.
+Our findings show a promising step towards modeling human context and provide valuable insights for the NLP community to think of additional strategies for improving models with task-dependent human context in pre-training.
 Limitations
@@
 However, similar effects are likely to hold for other languages, and should be evaluated in future work.
+Acknowledgments
+This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #585968.
+The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
 Appendix A Appendix
@@
 Pre-training takes approx 14 hours for 1 epoch and fine-tuning takes approx 1-4 hours depending on the task.
-Age bucket
-#Users
-HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
-<18
-503
-0.627
-0.644
-0.618
-18-21
-560
-0.557
-0.608
-0.592
-21-30
-563
-0.715
-0.741
-0.738
-30-45
-249
-0.594
-0.669
-0.667
-45+
-68
-0.567
-0.546
-0.599
+Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
+<18 503 0.627 0.644 0.618
+18-21 560 0.557 0.608 0.592
+21-30 563 0.715 0.741 0.738
+30-45 249 0.594 0.669 0.667
+45+ 68 0.567 0.546 0.599
 Table 6: Disattenuated pearson $r$ for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row.
-Age bucket
-#Users
-HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
-<18
-1113
-4.07
-2.52
-2.82
-18-21
-1387
-6.52
-4.00
-3.89
-21-30
-1557
-17.82
-12.64
-13.11
-30-45
-695
-48.59
-39.79
-40.43
-45+
-248
-114.92
-121.66
-134.72
+Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
+<18 1113 4.07 2.52 2.82
+18-21 1387 6.52 4.00 3.89
+21-30 1557 17.82 12.64 13.11
+30-45 695 48.59 39.79 40.43
+45+ 248 114.92 121.66 134.72
 Table 7: Mean squared error for age over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).
-Age bucket
-#Users
-HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
-<18
-503
-0.423
-0.410
-0.429
-18-21
-560
-0.496
-0.487
-0.506
-21-30
-563
-0.429
-0.380
-0.381
-30-45
-249
-0.578
-0.489
-0.489
-45+
-68
-0.584
-0.501
-0.467
+Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
+<18 503 0.423 0.410 0.429
+18-21 560 0.496 0.487 0.506
+21-30 563 0.429 0.380 0.381
+30-45 249 0.578 0.489 0.489
+45+ 68 0.584 0.501 0.467
 Table 8: Mean squared error for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).

### v3

@@
 Abstract
-Incorporating human context into language models is the next frontier for human-centered natural language processing.
-Currently, two pre-training methods exist: group-wise attributes (e.g., over-45-year-olds) or individual traits.
-Group attributes are coarse — not all 45-year-olds write the same way — while modeling individual traits allows for a more personalized representation, but requires more complex modeling and data.
-So far, it is unclear which pre-training approach benefits what tasks.
-We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on 5 user- and document-level tasks.
-We find that pre-training with both group and individual features significantly improves the two user-level regression tasks like age estimation and personality assessment.
-Pre-training on individual users significantly improves the three document-level classification tasks like stance and topic detection.
-It even does well for downstream tasks without historical user data.
-Our results suggest both approaches have specific use cases, opening new avenues for human-centered language modeling.
+Pre-trained language models consider the context of neighboring words and documents but lack any author context of the human generating the text.
+However, language depends on the author’s states, traits, social, situational, and environmental attributes, collectively referred to as human context (Soni et al., 2024).
+Human-centered natural language processing requires incorporating human context into language models.
+Currently, two methods exist: pre-training with 1) group-wise attributes (e.g., over-45-year-olds) or 2) individual traits.
+Group attributes are simple but coarse — not all 45-year-olds write the same way — while individual traits allow for more personalized representations, but require more complex modeling and data.
+It is unclear which approach benefits what tasks.
+We compare pre-training models with human context via 1) group attributes, 2) individual users, and 3) a combined approach on five user- and document-level tasks.
+Our results show that there is no best approach, but that human-centered language modeling holds avenues for different methods.
 Comparing Pre-trained Human Language Models:
@@
 1 Introduction
-Language varies between people.
-To capture this notion, two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language.
-The first approach focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group.
+Language is a fundamental form of human expression that varies between people.
+Pre-trained Language Models (PLMs) account for the textual context of neighboring words and documents but lack the human context of the author “generating" the language.
+However, language is highly dependent on the human context Soni et al.
+(2024), i.e., an author’s changing states Fleeson (2001); Mehl and Pennebaker (2003), traits, social, situational, and environmental attributes.
+For example, a person’s language differs when hiking (situation/environment) versus when feeling dejected (state) over a breakup (situation).
+It is essential to model the additional human context to better understand human language with PLMs Soni et al.
+(2024).
+Two strands of human-centered Natural Language Processing (NLP) emerged to model the people behind the language.
+The first focuses on the group context, building on the sociolinguistic notion of specific socio-demographic attributes influencing the language of a particular group.
 These socio-demographic attributes include age, gender (Volkova et al., 2013; Hovy, 2015), location (Kulkarni et al., 2016; Garimella et al., 2017), personality (Schwartz et al., 2013; Lynn et al., 2017), and more.
-The second approach focuses on building personalized language models (PLMs) that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
+The second approach focuses on building personalized language models that target individualistic contexts (King and Cook, 2020; Delasalles et al., 2019), and latent attributes inferred from an individual’s historical language (Matero et al., 2021; Soni et al., 2022) to better model the user.
 Figure 1: Pre-training a language model with no human context, with socio-demographic group attribute, with individual traits, and with both group and individual traits.
-While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2023): People are not defined by their group membership alone Orlikowski et al.
-(2023), but individual traits might not generalize.
-We compare the downstream performance of models pre-trained with the two approaches as well as their combination.
+While these two strands have advanced human-centered NLP, we still do not understand their relative strengths, complementarity, and impact on different tasks (Soni et al., 2024).
+People are not defined by their group memberships alone Orlikowski et al.
+(2023), and individual traits might not be generalizable enough across groups.
+Further, cross-cultural psychology research (Hofstede and Bond, 1984) notes the importance of both individualism and collectivism and Soni et al.
+suggest a flexible interplay of these concepts when including human context in PLMs.
+We might expect models informed by both group and individual features to perform better, but no data exists on this.
+In this work, we empirically evaluate these hypotheses and compare the effects of including different types of human context in PLMs (i.e., groups context (collectivism), individualistic aspects (individualism), and a combination of both) on specific tasks.
 We answer the following broader research questions:
-(RQ1): How can we incorporate group and/or individual human context into pre-training?
-(RQ2): How does pre-training with different human contexts differ in terms of downstream performance for different tasks?
-We use the monolingual socio-demographically adapted model from Hung et al.
-(2023) and the HaRT model from Soni et al.
-(2022) for group and individual human contexts, respectively.
-For a model trained with both individual and group human context, we use a multi-task learning PLM based on HaRT to create a model, GRIT, in two variants: GRITage is adapted to the authors’ age, and GRITope to their inferred personality trait (openness).
-We test all systems on five user- and document-level tasks.
-Note that because we focus on empirically comparing pre-training group and individual traits in PLMs, we cannot compare to large language models like GPT4, which do not support stratification to either attribute via pre-training.
-Recent studies have explored methodologies like user-adapters (Zhong et al., 2021), and user-centric prompting (Li et al., 2023) to include human context into models.
-In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts.
-We find that PLM pre-trained on individuals and group features enhances 2 user-level regression tasks: age estimation and personality assessment from multiple user documents.
-These user-level tasks focus on individual people.
-Our findings, that they are best modeled as a combination of their group attributes and unique characteristics, conforms with the notion that a person is a mix of both.
-Document-level categorization tasks, like stance detection, are more personal than group based.
-A PLM pre-trained on individual human context alone improves 3 document-level classification tasks.
+(RQ1): How can we incorporate group and individual human context into pre-training?
+(RQ2): How does pre-training with different human contexts affect downstream performance for different tasks?
+Recent works trained PLMs with socio-demographic group context (Hung et al., 2023), and individual context (Soni et al., 2022).
+However, no empirical studies compare the impacts of different types of human contexts included in language modeling.
+We compare the downstream tasks’ performance of models from these works with that of a new PLM trained with group and individual contexts.
+We test all systems on five downstream tasks from these works to ensure a variety of tasks and prediction properties at three levels: 1) user level, with historical language from authors (age estimation and personality assessment), 2) document-level, with historical language from some authors (stance detection), 3) document-level, without historical language from authors (topic detection and age category classification).
+Note that because we focus on empirically comparing pre-training with different types of human context, we cannot compare to the larger language models like GPT4, which are not yet pre-trained/trainable with human context but are envisioned to become large human language models in the future (Soni et al., 2024).
+Recent studies have explored methodologies like user-adapters (Zhong et al., 2021) and user-centric prompting (Li et al., 2023) to include human context into the larger language models.
+In contrast, we focus specifically on comparing the impact of pre-training LMs with different human contexts, as Soni et al.
+(2024) argue that pre-training will allow for modeling a richer human context by explicitly handling the multi-level structure of documents connected to people, as opposed to specific and limited benefits of user-centric prompting and fine-tuning (Salemi et al., 2023; Choi et al., 2023).
+PLMs pre-trained on individual and group features enhance performance on two user-level regression tasks that use multiple user documents: age estimation and personality assessment.
+In contrast, PLMs pre-trained on individual human context alone improve performance on document-level classification tasks like stance and topic detection.
+Our findings suggest user-level tasks focusing on individual people are best modeled as a mix of both group attributes and unique characteristics.
+However, document-level tasks that are more personal, like stance detection, are best addressed by modeling the individual context alone.
+Adding group attributes degrades performance.
 By their very nature, models of this kind touch upon sensitive user information.
-For this reason, we take a responsible release strategy, making only the code for the comparisons publicly available and the exact splits of the TrustPilot and Stance datasets used.
-We build on top of the publicly available code for HaRT and Hung et al.
+For this reason, we adopt a responsible release strategy, making only the code for the comparisons publicly available, along with the exact splits of the TrustPilot and Stance datasets used [^1] https://github.com/soni-n/HumanContextLanguageModels_Comparison.
+We build on top of the publicly available code from Soni et al.
+(2022) and Hung et al.
 (2023).
-We acquired the model and data in a secure manner from the authors of Soni et al.
-(2022) and TrustPilot data splits from the authors of Hung et al.
+We acquired the models and data from the authors of the respective works in a secure manner.
+For more information about the models and data, see Sections 4 and 5.
+For a discussion of the ethical implications of the models and data, see the Ethical Considerations section.
+Contributions.
+(1) We empirically compare three pre-training strategies for language models with human context: individual traits, group socio-demographic features, and a combination of both.
+(2) We evaluate each pre-training strategy on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category classification).
+(3) We find that the two user-level regression tasks perform better with PLMs pre-trained with individual and group human contexts.
+Conversely, the three single document-level tasks perform better with PLMs pre-trained with individual context alone.
+Further, pre-training with group and individual contexts reduces performance for document-level tasks.
+2 Related Work
+Socio-demographic and latent human factors.
+Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models.
+Initial studies show benefits of grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups, and improving text classification tasks like sentiment analysis.
+Additionally, adapting to socio-demographic user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) have been effective to enhance the performance in tasks like sarcasm detection, and toxic language detection.
+Some studies go beyond modeling explicit groups, to learn individual representations latently Jaech and Ostendorf (2018); Delasalles et al.
+(2019) or via historical language Matero et al.
 (2021).
-For more information about the model and data, see Sections 3 and 4.
-For a discussion of the ethical implications of the model and data, see Ethical Considerations section.
-Contributions.
-(1) We compare pre-training several language models with individual users, group socio-demographic features, and both traits.
-(2) We evaluate the three specific pre-training strategies on five downstream tasks: two multi-document user-level regression (personality-openness evaluation and age estimation) and three single document-level classification tasks (stance detection, topic detection, and age category prediction).
-(3) We find that the two user-level regression tasks perform better when pre-trained with both strategies (GRIT) and the three single document-level tasks perform better on fine-tuning a language model pre-trained with individual context alone (HaRT).
+Pre-traing with human context.
+With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks.
+More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context and evaluate on personalized and social tasks, resulting in mediocre performance (Salemi et al., 2023; Choi et al., 2023).
+However, few studies have explored including human context within the pre-training regime of LMs.
+Hung et al.
+(2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup.
+Additionally, Soni et al.
+(2022) pre-train a LM with individual human context derived from user’s historical language.
+Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
 3 Integrating Human Context in PLMs
-For our comparison, we use three systems representing the paradigms of pre-training with human context (Figure 1).
-We want to tease apart the contributions of 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts.
-As noted earlier, we focus on recent approaches for pre-training language models with the additional human context.
+For our comparison, we use three systems representing the three paradigms of pre-training with human context (Figure 1).
+We want to tease apart the contributions of different types of human context: 1) grouping people, 2) modeling individual users, and 3) modeling both group and individual human contexts.
+As noted earlier, we focus on recent approaches for pre-training language models with additional human context.
 Pre-training with group context.
-We build on Hung et al.
-(2023), a recent model to explore demographic adaptation in transformer-based PLMs.
-They use bidirectional auto-encoder-based PLMs to inject demographic knowledge in a multi-task learning setup where they also train masked language modeling (MLM) and classify the gender or age of an author.
-They use the multilingual reviews dataset with demographic labels from Trustpilot [^1] https://www.trustpilot.com/ Hovy (2015).
-They evaluate multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection.
-We use the US-English subset of the Trustpilot data for topic detection (TD) under the age categories and for age attribute classification (AC) (more details in section 4).
-We compare to the results from Hung et al.
-(2023) with monolingual BERT on a US-English dataset with out-of-domain demographic (age) specialization.
-Out-of-domain data is the Blogs authorship corpus Schler et al.
-(2006), and in-domain means Trustpilot corpus.
-To be consistent and fair in comparing with other human context pre-training paradigms, we choose the monolingual model and eliminate the confounds from domain specialization.
+We build on a model from Hung et al.
+(2023) that explores demographic adaptation in transformer-based PLMs.
+It is a bidirectional auto-encoder-based PLM injecting demographic knowledge in a multi-task learning setup where they train for masked language modeling (MLM) and classify the gender or age of an author.
+They use the Trustpilot dataset  [^2] https://www.trustpilot.com/ of multilingual reviews with demographic labels Hovy (2015), and evaluate on multiple text classification tasks, including demographic attribute classification, sentiment analysis, and topic detection.
+For our comparison study, we use the US-English subset of the Trustpilot data for two tasks: topic detection (TD) across two age categories, and age attribute classification (AC) (more details in section 5).
+Additionally, we use a monolingual BERT pre-trained with age specialization on the Blogs authorship corpus Schler et al.
+(2006).
+This choice allows us to eliminate a domain influence (i.e., Trustpilot reviews), given that the other PLMs under comparison lack this specialization.
 Pre-training with individual human context.
 Soni et al.
-(2022) introduced human language modeling (HuLM) in pre-trained LMs.
-I.e., pre-training a regular language modeling but including a dynamic individual human context vector derived from the authors’ historical texts.
-This vector captures the human states in which the text was generated.
+(2022) introduced human language modeling (HuLM) in PLMs, which is regular language modeling given an additional context of the individual generating the language.
+This additional context is a dynamic vector derived from the authors’ historical texts motivated by the idea of capturing the changing human states expressed in language.
 It also adds coherence to texts generated by the same author.
-They introduce a Human-aware Recurrent Transformer (HaRT), an autoregressive PLM for human language modeling.
-They evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks.
-We use the results from HaRT on the user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
+They introduce Human-aware Recurrent Transformer (HaRT), an autoregressive PLM to evaluate the effect of individual human context on language modeling and multiple user-level and document-level downstream tasks.
+We build on HaRT and use two user-level tasks, age estimation and personality (openness) assessment, and on a document-level task, stance detection, for our comparisons study.
 Pre-training with both group and individual human context.
-We train a PLM to integrate both individual and group human context in language modeling pre-training.
-We extend HaRT by predicting an additional human group attribute in a dynamic muti-task learning setup similar to Hung et al..
-To facilitate comparison, we combine the pre-training inducing the individual human context through the author’s language and inject the group context by predicting a group attribute of the author.
-We choose downstream tasks from both prior works (Soni et al., 2022; Hung et al., 2023) to compare our three language models pre-trained with different levels of human contexts: two user-level regression tasks and three single document-level classification task.
+We train a PLM to integrate both individual and group human context by introducing a multi-task learning setup into HaRT that incorporates group features.
+This approach facilitates training a PLM with both group and individual context.
+We evaluate the model on two multi-document user-level regression tasks: age estimation and personality assessment, and three single document-level classification tasks: stance detection, topic detection, and age group classification.
+Importantly, the only difference in this multi-task learning setup compared to HaRT is the inclusion of a demographic attribute prediction during pre-training, similar to how Hung et al.
+(2023) adapted traditional PLMs for group context.
 4 Models
@@
 Soni et al.
-(2022) introduced HaRT to incorporate individual human context into PLMs.
-They use a 12-layered autoregressive GPT-2 based architecture with a modified self-attention computation at layer 2.
+(2022) use a 12-layered autoregressive GPT-2 based architecture with a modified self-attention computation at layer 2.
 This modification to the query vector now includes the individual human context via a dynamic user-state vector.
@@
 (2023) explore socio-demographic adapted BERT models to inject group human context into PLMs.
-We use the names BERTDS and BERTage-MLM to denote their demographic (age) specialization using the multi-task learning setup and demographic adaptation with masked language modeling respectively.
+We use two models: one specialized for age (demographic attribute) under the multi-task learning setup, and the other adapted to the age corpora through standard masked language modeling.
+We denote these as BERTDS and BERTage-MLM, respectively.
 Multi-Task Learning.
 Hung et al.
-(2023) train for domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$).
+(2023) train for both domain adaptation using the masked language modeling ($L_{mlm}$) loss and for classifying demographic category using the binary cross-entropy loss ($L_{dem}$).
+Both losses must be combined to simultaneously learn multiple objectives.
 To account for the homoscedastic uncertainty Kendall et al.
-(2018) of losses, they adopt a dynamic MTL objective for training with group human context.
-Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions.
-This approach accounts for the different scales across multiple losses.
-$\tilde{L_{t}}=\frac{1}{2\sigma_{mse}^{2}}L_{t}+\log\sigma_{t}$
+(2018) of both losses, they adopt a dynamic multi-task learning (MTL) objective for training with group human context.
+Homoscedastic uncertainty is a task-dependent weighting to derive a multi-task loss function that can optimally learn the weights and balance the impact of multiple loss functions and their different scales.
+The tasks are dynamically weighted using the variance of the task-specific loss ($\sigma_{t}^{2}$) over training instances of the task $t\in\{mlm,dem\}$:
+$\tilde{L_{t}}=\frac{1}{2\sigma_{t}^{2}}L_{t}+\log\sigma_{t}$
 Hung et al. minimize the sum of both the uncertainty adjusted losses: $\tilde{L}_{mlm}+\tilde{L}_{dem}$.
@@
 GRIT.
-GRoup and Individual HaRT is a LM pre-trained for the HuLM task and a (continuous) socio-demographic attribute prediction regression task in a multi-task learning setup.
-The PLM uses the user-state vectors to predict the socio-demographic attribute of the user.
+We train HaRT under a multi-task learning setup for both the individual context — through the HuLM pre-training task (see Section 4.1) — and the group features — via a regression task to predict a (continuous) socio-demographic attribute of the author.
+We call the model as GRoup and Individual HaRT (GRIT).
+The model uses the user-state vectors (see Section 4.1) to predict the socio-demographic attribute of the author:
 $Pr(attribute|\overline{\mathbf{U}})$
-Because of compute limitations, we chunk a user’s language history into blocks and process them in a single forward pass.
+We chunk a user’s language history into blocks and process them in a single forward pass.
 Each block of text from a user results in a user-state vector.
@@
 This representation is layer-normed and linearly transformed before making a continuous-valued prediction for the specific attribute.
-We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope), respectively.
-The models train on a regression loss for the attribute prediction regression tasks using mean squared error ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$).
-Similar to Hung et al.
-(2023), we account for the homoscedastic uncertainty Kendall et al.
-(2018) of losses.
-We use the joint loss for a continuous and discrete output as derived in Kendall et al.
+We pre-train one model for the continuous attribute age (GRITage) and one for the continuous attribute personality type openness (GRITope).
+The models train on a regression loss for the attribute prediction regression tasks using mean squared error loss ($L_{mse}$), and a classification loss for the HuLM task using cross-entropy loss ($L_{ce}$).
+We must combine both losses to jointly learn the two objectives and account for the homoscedastic uncertainty Kendall et al.
+(2018) of the losses.
+Since we combine a regression and a classification loss, we train the model to learn to balance the loss for a continuous and discrete output as derived in Kendall et al.
 (2018) and compute our joint objective as follows:
 $\frac{1}{\sigma_{ce}^{2}}L_{ce}+\frac{1}{2\sigma_{mse}^{2}}L_{mse}+\log\sigma_{ce}+\log\sigma_{mse}$
-where, $\sigma_{ce}$ and $\sigma_{mse}$ are the variances of the task-specific losses.
-In practice, we use log variance for numerical stability and use the adjusted loss calculation as follows:
-$\exp^{-\eta_{ce}}L_{ce}+\eta_{ce}+\frac{1}{2}(\exp^{-\eta_{mse}}L_{mse}+\eta_{mse})$
-where $\eta_{x}=\log\sigma_{x}^{2}$.
+where, $\sigma_{ce}^{2}$ and $\sigma_{mse}^{2}$ are the variances of the task-specific losses over the training instances of the respective tasks.
+To add numerical stability, we adjust the loss calculation to use log of the variance:
+$\exp^{-\eta_{ce}}L_{ce}+\frac{1}{2}(\exp^{-\eta_{mse}}L_{mse}+\eta_{ce}+\eta_{mse})$
+where $\eta_{x}=\log\sigma_{x}^{2}$ for $x\in\{mse,ce\}$.
 We let $\sigma_{ce}$ and $\sigma_{mse}$ be learnable parameters for the model.
-We do not halve the log-term of the cross-entropy loss since we found it to perform better with our multi-task learning experiments.
-Pre-training data.
-We use a subset of the pre-training data for HaRT, consisting of the demographics and personality information.
-This subset contains the Facebook posts from Park et al.
-(2015) as used by Soni et al..
-Our dataset is consistent with the inclusion criteria for HaRT to ensure moderate language history for each user: we include English posts from users with at least 50 total posts and at least 1000 words.
-This dataset consists of just over 63,000 unique users, which we split into a training dataset consisting of messages from 56,930 users, a development dataset that consists of messages from 1836 users that were not part of the training set, and a test set of messages from a separate set of 4438 users that are neither in training nor the development set.
-To evaluate the human attribute prediction in GRITope, we use a subset of the test set consisting of messages from 1745 users to accommodate for questionnaire reliability.
-We use the Facebook posts for the HuLM task and the demographic and personality scores of consenting Facebook users (Kosinski et al., 2013) for the human attribute prediction task.
-Training.
-We use HaRT’s pre-trained weights as the base weights for GRIT and randomly initialize the newly introduced weights for human attribute prediction.
-GRIT is trained on our pre-training dataset using the 5e-5 learning rate after experimenting with a few learning rates, including that used for HaRT’s pre-training.
-Following HaRT, and due to computing limitations, each training instance is capped to 8 blocks of 1024 tokens each, with train batch size as 1 per device and evaluation batch size as 20 per device, trained over 2 GPUs for eight epochs.
-We explored multiple joint losses before resorting to the homoscedastic loss computation.
-Since HaRT caps to 4 train blocks for user-level downstream tasks, we also pre-train GRITage and GRITope with four training blocks.
+In practice, we do not halve the $\eta_{ce}$ term in the above equation since we found it to perform better with our multi-task learning experiments.
+Pre-training Data and Training.
+We use the same Facebook posts dataset (Park et al., 2015) and training, validation, and test splits as those used by Soni et al.
+(2022).
+For both GRITage and GRITope, we use the demographic and personality scores, respectively, obtained from consenting Facebook users (Kosinski et al., 2013).
+This data is identical to that used by HaRT for the age estimation and personality assessment tasks.
+During training, we use a learning rate of 5e-5 in the multi-tasking training setup, employing the homoscedastic loss computation method described earlier.
+Following the experimental settings for HaRT, each training instance is capped to 4 blocks of 1024 tokens each.
+We use a train batch size of 1 per device and an evaluation batch size of 20 per device, trained over 2 GPUs for 8 epochs.
+Further details can be found in Appendix A.1.
 4.4 Fine-Tuning
-We fine-tune GRIT and HaRT for downstream document-level tasks.
-Each downstream task has a separate fine-tuned model that is initialized with the respective model’s pre-trained parameters and trained using the respective downstream task labels and an appropriate loss function.
-We also use the historical language of a user where available for any of the downstream tasks.
-We use the last predicted token’s representation to predict the label in document-level classification tasks.
-We experimented with fine-tuning GRIT for user-level regression tasks in multiple ways, including 1) similar to HaRT, by using the averaged user-state vectors from GRIT, 2) same as previous but fine-tuning only the history module, attribute prediction module, and the downstream task head, 3) freezing all the parameters of GRIT and fine-tuning the human attribute prediction module alone.
-However, we found continued training as described in section 3.5 to perform best.
-We used the Optuna framework Akiba et al. (2019) for hyperparameter search, closely following the experimental settings in Soni et al..
+We utilize the results of fine-tuned BERTDS and BERTage-MLM from Hung et al.
+(2023), as well as fine-tuned HaRT models from Soni et al.
+(2022) where available.
+We fine-tune both GRIT models for all downstream tasks, and HaRT for 2 document-level tasks.
+Additionally, we use the Optuna framework Akiba et al.
+(2019) for hyperparameter search, closely following the experimental settings in Soni et al.
+(2022).
+Details can be found in Appendix A.2.
 4.5 Transfer Learning
-We experiment with continuing pre-training GRIT models for each group attribute.
-To this end, we pre-train GRITage capped to 4 training blocks and use this pre-trained model to continue MTL with the HuLM task and predict personality (openness).
-We do the same for GRITope and continued MTL with predicting age.
+We experiment with fine-tuning GRITage in a multi-task learning setup for both the HuLM task and predicting personality (openness).
+Similarly, we fine-tune GRITope to predict age while also training for the HuLM task.
+We observe that this form of transfer learning yields the best performance for the user-level regression tasks (refer to Section 6.1).
 5 Experiments
-Our research goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined.
-To this end, we evaluate performances of the models defined in Section 3 on two user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification.
+Our study’s goal is to compare the downstream performance of models pre-trained with human contexts in three forms: socio-demographic group factors, individual traits, and combined.
+To this end, we evaluate performances of the models defined in Section 4 on two multi-document user-level regression tasks: predicting age and a personality score (openness), and on three single document-level classification tasks: stance detection, topic detection, and age classification.
 We also compare against GPT-2HLC from Soni et al.
 (2022) as a PLM adapted to the social media domain but devoid of human context.
+All experiments were run using Optuna trials (Akiba et al., 2019) to search for the best hyperparameters and reduce the effects of randomness.
+More details are provided in Appendix A.2
 5.1 User Level Regression Tasks
-We compare GRIT, HaRT, and GPT-2HLC on age estimation and personality (openness) assessment Kosinski et al.
-(2013); Park et al.
-(2015).
-These social scientific tasks require predicting continuous outcomes for a user given multiple documents written by them.
+We consider two user-level social scientific tasks: age estimation, and personality (openness) assessment, which require predicting continuous outcomes (real-valued age, or openness score) for a user given multiple documents written by them.
 We use the same data splits as used by Soni et al.
-(2022) for evaluating HaRT and GPT-2HLC
-We use the pre-trained GRITage and GRITope directly to evaluate on the test sets for age estimation and personality assessment, respectively.
-We further evaluate these models on the test sets for personality assessment and age estimation, respectively, after continuing training for these tasks, as described in section 3.5.
-We use the results from Soni et al.
-(2022) for HaRT and GPT-2HLC which are directly comparable to GRIT models trained on the same data splits and metrics.
-Soni et al.
-fine-tuned the recurrence module of the pre-trained HaRT model for the tasks of age estimation and personality assessment using the average of user-states from non-padded blocks of texts from an author, resulting in two fine-tuned models.
-Similarly, they fine-tune the last two layers of the pre-trained GPT-2HLC model for these tasks.
-Since GPT-2HLC can not handle all text from a user in one pass, they average the predictions across all user messages corresponding to the same label for each message.
+(2022) for our comparison study.
+Since GRITage is pre-trained using age estimation as one of the tasks, we use directly evaluate it on the held-out test set.
+This allows for direct comparison with HaRT fine-tuned for the age estimation task.
+Furthermore, we can potentially attribute performance differences to the training with combined group and individual context, as GRITage incorporates the group feature into HaRT’s architecture.
+Similarly, GRITope is evaluated on the held-out test set for personality assessment.
+Moreover, we evaluate GRITage and GRITope for the tasks of personality assessment and age estimation, respectively, using the transfer learning mechanism described in Section 4.5.
+We report and compare pearson $r$ for age estimation and disattenuated pearson $r$ for personality assessment.
 5.2 Document-Level Classification Tasks
 We compare different models for stance detection vs.
-topic detection and age classification.
-These tasks classify a single input document (tweet message or a review) a user writes into label categories.
-For stance detection, we also use the historical messages of a user where available, as in Soni et al.
+topic detection and age classification tasks.
+These tasks classify a single input document (tweet message or a review) that a user writes into label categories.
+For stance detection, we also utilize the historical messages of a user where available, as in Soni et al.
 (2022).
-We do not have the user information or any user historical language available for the other two tasks, so we evaluate on the single document input.
+However, we do not have the user information or any user historical language available for the other two tasks, so we evaluate solely based on the single document input.
 All models process the input document(s) and feed the layer-normed last non-padded token representation to the classification layer to classify the document into label categories.
-Only GRIT and HaRT incorporate user information and the historical language (where available).
-The other two models can only use the input document without a hierarchical structure to make the predictions.
-We compare with the results from Soni et al.
+Only GRIT and HaRT incorporate user information and the historical language available for the stance detection task.
+However, GPT-2HLC, and both BERTDS and BERTage-MLM lack this hierarchical structure and can only use the input document without access to historical data for making predictions.
+We compare the results from Soni et al.
 (2022) and Hung et al.
@@
 Given a single annotated tweet, this task predicts a user’s stance as in favor of, against, or neutral towards one of the five targets: atheism, climate change as a real concern, feminism, Hillary Clinton, and legalization of abortion.
-We fine-tune the models under comparison for each target separately.
-We report average of weighted F1 scores with three labels across all five targets.
-We use Soni et al.
-(2022)’s train/dev/test split over SemEval 2016 dataset (Mohammad et al., 2016).
+We fine-tune GRITage and GRITope for each target separately, and use the results from Soni et al.
+(2022) for GPT-2HLC and HaRT.
+We report the average of weighted F1 scores [^3] We use this metric to maintain consistency with previous works under comparison (Soni et al., 2022; Hung et al., 2023).
+with three labels across all five targets.
+We use the train/dev/test split provided by Soni et al.
+(2022) over the SemEval 2016 dataset (Mohammad et al., 2016).
 HaRT and GRIT models maintain the temporal accuracy by using only the messages posted earlier than the labeled messages from the extended dataset (Lynn et al., 2019) as a user’s historical language.
-We compare the results of fine-tuned GPT-2HLC, HaRT (Soni et al., 2022), and fine-tuned GRITage and GRITope.
 Topic Detection
 We use the US subset of the TrustPilot reviews dataset Hovy (2015) from two age groups: below 35 or above 45  [^4] As suggested by Hovy (2015), this split of the age ranges results in roughly equally-sized data sets and is non-contiguous, avoiding fuzzy boundaries..
-Given a single review, this task predicts the review topics from five categories: Flights, Online marketplace, Fitness & Nutrition, Electronics, and Hotels.
-We use the same train, development, and test set splits as Hung et al.
-(2023) to eliminate any skew in the demographically-conditioned label distribution.
-We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
+Given a single review, the task is to predict the review topics from five categories: Flights, Online marketplace, Fitness & Nutrition, Electronics, and Hotels.
+To maintain consistency, we adopt the same train, development, and test set splits as Hung et al.
+(2023) to ensure a stratified demographically-conditioned label distribution.
+We fine-tune GPT-2HLC, HaRT, GRITage, and GRITope using these data splits to predict the topic for a given review, and report macro-F1 scoresfootnotemark: .
+We also compare to results from BERTage-MLM and BERTDS (Hung et al., 2023).
 Demographic Attribute Classification
@@
 Age categories are equally represented in each set.
-We report and compare macro-F1 scores from BERTage-MLM and BERTDS (Hung et al., 2023) with fine-tuned GPT-2HLC, HaRT, GRITage and GRITope.
-4.3 Human Language Modeling
-To compare the effects of individual and group factors on language modeling performance, we evaluate on the test set from the pre-trained data splits.
-We report and compare perplexity scores from the pre-trained GPT-2 (GPT-2frozen), GPT-2HLC, HaRT, GRITage and GRITope for the human language modeling task.
-5 Results and Discussion
-We report results for all the tasks here, discussing their respective impacts from pre-training LMs with individual human context, group context, and both individual and group context.
-5.1 Comparisons Study
-User-Level Regression Tasks.
-Table 1 shows the two user-level regression task results.
-We find that just the pre-trained GRIT models for each task perform better than the fine-tuned HaRT model, i.e., pre-trained GRITage better estimates age, and GRITope better assesses personality.
-Additionally, comparing the transfer learning (Section 3.5) results of GRITage for openness and GRITope for age to the fine-tuned HaRT and GPT-2HLC models, we find the GRIT models to fare better.
-Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it is still devoid of human context.
-HaRT adds pre-training with the individual human context, and GRIT adds pre-training with both group and individual human contexts (Figure 1).
-As Table 1 shows, there are gains going from GPT-2HLC (no human context) to HaRT (individual human context) and further to GRIT (individual + group human context).
-This indicates that pre-training PLMs with individual and group human context may benefit multi-document user-level regression tasks like the ones we considered.
+We fine-tune GPT-2HLC, HaRT, GRITage and GRITope using the provided splits to predict if the review is written by someone below 35 years or above 45 years, and report macro-F1 scoresfootnotemark: .
+We also compare to results from BERTage-MLM and BERTDS (Hung et al., 2023).
 Model Human Context Age ($r$) OPE ($r_{dis}$)
@@
 All other results show statistical significance $p<0.05$ using paired t-test.
+5.3 Human Language Modeling
+To compare the effects of individual and group factors on language modeling performance, we evaluate on the test set from the pre-trained data splits.
+We report and compare perplexity scores from the pre-trained GPT-2 (GPT-2frozen), GPT-2HLC, HaRT, GRITage and GRITope for the human language modeling task.
+Model Human Context Stance (F1wtd) TD (<35) (F1mac) TD (>45) (F1mac) AC (F1mac)
+GPT-2HLC None 68.60 69.77 65.43 63.93
+BERTage-MLM Group - 68.40 64.60 61.90
+BERTDS Group - 69.30 65.00 64.10
+HaRT Individual 71.10 69.84 65.65 64.33
+GRITage Ind + Grp 70.82 69.21 64.52 62.56
+GRITope Ind + Grp 70.07 66.53 64.84 61.18
+Table 2: Weighted F1 for stance detection, macro-F1 for topic detection (TD), and age classification (AC) on TrustPilot reviews.
+Pre-training with individual context appear to benefit all tasks.
+Bold $=$ best in column; McNemar’s test comparing classifiers does not show statistical significance between the best performing model (HaRT) and the best baseline with no individual context (GPT-2HLC).
+6 Results and Discussion
+We report results for all the tasks here, discussing their respective impacts from pre-training LMs with individual human context, group context, and both individual and group context.
+6.1 Comparisons Study
+User-Level Regression Tasks.
+Table 1 shows the results of the two user-level regression tasks.
+We find that GRIT models outperform others for both age estimation and personality assessment tasks.
+Additionally, upon comparing the transfer learning (Section 4.5) outcomes of GRITage for openness and GRITope for age to those of the HaRT and GPT-2HLC models, we consistently observe superior performance with the GRIT models, further substantiating their efficacy.
+Note that while GPT-2HLC is a PLM that is adapted to the social-media domain, it lacks human context.
+HaRT incorporates individual human context in pre-training, and GRIT extends this by integrating both group and individual human contexts in pre-training (Figure 1).
+As shown in Table 1, there are gains observed from GPT-2HLC (no human context) to HaRT (individual human context), and further to GRIT (individual + group human context).
+This suggests that pre-training PLMs with individual and group human context can benefit multi-document user-level regression tasks, such as those we considered.
+Importantly, the only difference between HaRT and GRIT models lies in the integration of the demographic attribute prediction (group context).
+Both models are pre-trained and evaluated on precisely the same data, allowing performance differences to be attributed to the additional group context combined with individualistic human context.
 Document-Level Classification Tasks.
-Model Human Context Stance (F1wtd) TD (<35) (F1mac) TD (>45) (F1mac) AC (F1mac)
-GPT-2HLC None 68.6 69.8 65.4 63.9
-BERTage-MLM Group - 68.4 64.6 $61.9$
-BERTDS Group - 69.3 65.0 $64.1$
-HaRT Individual $\textbf{71.1}^{*}$ $\textbf{69.8}^{*}$ 65.6 $\textbf{64.3}^{*}$
-GRITage Ind + Grp 70.8 69.2 64.5 62.7
-GRITope Ind + Grp 70.1 66.5 64.8 61.2
-[…]
-Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute classification (AC).
-We see that task fine-tuned HaRT (individual human context) models performs better on all tasks..
-HaRT models inherently have an additional context of the individual user and do not treat all inputs as if written by the same user.
-The tasks considered here relate more to personal opinions and preferences, rather than group-level ones.
-HaRT model is well-suited for incorporating such personalization due to its pre-training with individual human context.
-Additionally, the results indicate that the models pre-trained with group context (BERTDS) fare well in the group-based tasks of topic detection and age classification.
-Whereas, the models pre-trained on both individual and group human context (GRIT) appears to bring in noise for the group-based and personal stance detection tasks resulting in a slightly worse performance.
-Further, it is important to note that the individual human context (HaRT) derived from the historical tweets from the users in the stance detection dataset provides for a richer human context.
-Although, the performance is not greatly hurt even if historical language is not available for certain tasks (TD and AC).
-Perplexity.
-We also compare the language modeling capability of the various models.
-Table 3 reports perplexity on the test set of 4438 users.
-The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, HaRT (individual human context) models perform best while GRIT (individual + group human context) models result in a slightly lower perplexity than HaRT.
-GRIT models are pre-trained under a multi-task learning setup which most likely is hurting the individual task performances, thus resulting in a slight dip in the perplexity compared to HaRT.
-Further, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al.
-(2022).
+Table 2 shows the results for the 3 document-level classification tasks: stance detection, topic detection (TD) for 2 age groups (<35 and >45), and demographic attribute (age) group classification (AC).
+We see that task fine-tuned HaRT (individual human context) models perform better on all tasks.
+HaRT models inherently include an additional context of the individual user and do not treat all inputs as if written by the same user.
+The considered stance detection task primarily relates to personal opinions and preferences, rather than group-level ones, making HaRT well-suited for incorporating such personalization due to its pre-training with individual human context.
+While a group context may also influence a person’s stance to some extent, empirical observations show that the combination of individual and group contexts negatively impacts performance.
+Additionally, models pre-trained with group context (BERTDS) perform well in group-based tasks such as topic detection and age classification.
+However, models pre-trained on both individual and group human context (GRIT) do not appear to enhance results in group-based, and personal stance detection tasks resulting in slightly worse performance.
+Further, it is important to note that the individual human context (HaRT) derived for some of the users using their historical tweets, where available, in the stance detection dataset provides a richer human context as we see greater gains in the performance of HaRT over GPT-2HLC.
+Conversely, when historical language is not available for certain datasets (topic detection and attribute classification), HaRT does not perform worse than GPT-2HLC and may even achieve marginal gains due to the inherent human context in the model.
+However, we leave the evaluation of the impact of historical language on human context for future work.
 Model Human Context Test ($ppl$)
@@
 Table 3: Comparing perplexity on language modeling for models trained with individual and group contexts.
-5.2 Error Analysis and Disparity
-We conduct an error analysis as a function of a socio-demographic group attribute.
-Specifically, we measure GRIT and HaRT for error disparity (Shah et al., 2020) – – a systematic difference in error as a function of demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015).
-We analyzed outcome and error disparity on age and openness prediction from both the models with individual context (HaRT) and combined individual plus group context (GRIT).
-First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set.
-Then, we compare the performance of our models across these buckets.
-Results from table 4 indicate that pre-training with individual and group context together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age.
-We see similar trends for assessing openness personality (Appendix Tables 6, 8).
-This suggests that the group attribute prediction maybe acting as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
-To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, that is averaged by the number of pairs (Shah et al., 2020).
-A lower averaged sum of differences implies lesser errors as a function of the age groups.
-Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
+Perplexity.
+We also compare the language modeling capability of the various models.
+Table 3 reports perplexity on the held-out test set.
+The frozen GPT-2 performs poorly compared to the social media domain adapted GPT-2HLC, while HaRT model with individual human context perform the best.
+In contrast, GRIT models with both individual and group human context exhibit a slightly lower perplexity than HaRT.
+An individual’s language is inherently personal, yet it can also be influenced by their group context to some extent, thereby affecting the perplexity results in language modeling tasks.
+However, GRIT models pre-trained with both individual and group context yield slightly worse perplexity measures.
+Additionally, we observe similar trends in perplexity gains from GPT-2HLC (no human context) to HaRT (individual context) or GRIT (individual plus group context) as also demonstrated in Soni et al.
+(2022).
 Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
@@
 Table 5: Mean error disparity for age estimation and openness personality assessment over five age buckets. Bold indicates best in column (lower is better).
-6 Related Work
-Much work in human-centered NLP has focused on identifying and evaluating inclusion of human context in our models.
-Initial studies experimented with grouping people by socio-demographic factors like age or gender (Volkova et al., 2013; Hovy, 2015) and geographical region (Bamman et al., 2014; Garimella et al., 2017) to capture the variation in language usage and meaning among different groups.
-These works improved sentiment analysis, polarity classification, and topic detection.
-Other researchers explored adapting to user factors (Lynn et al., 2017), social networks (Huang et al., 2014; Radfar et al., 2020), and social media attributes (Bamman and Smith, 2015) to improve downstream tasks like sarcasm detection, and toxic language detection.
-Some studies go beyond explicit groups and learning individual representations latently Jaech and Ostendorf (2018); Delasalles et al.
-(2019) or via historical language Matero et al.
-(2021).
-With respect to pre-trained LMs, recent studies have used adapter-based methodology (Li et al., 2021; Zhong et al., 2021) to include individual human contexts for downstream tasks.
-More recently, large language models have used user-centric prompting (Li et al., 2023) to include human context.
-However, few studies have explored including human context within the pre-training regime of LMs.
-Hung et al.
-(2023) generalize the task-specific EMPATH-BERT (Guda et al., 2021) to create a PLM injected with demographic group information using a dynamic multi-task learning setup.
-Further, Soni et al.
-(2022) pre-train a LM with individual human context derived from user language.
-Our study aims at comparing the impacts of pre-training LMs with individual, or group, or combined individual plus group human context.
+6.2 Error Analysis and Disparity
+We conduct an error analysis based on a socio-demographic group attribute (age groups), specifically focusing on age and openness prediction tasks.
+We measure the performance of GRIT and HaRT in terms of error disparity (Shah et al., 2020) — a systematic difference in error based on demographics as exemplified by the “Wall Street Journal Effect” Hovy and Søgaard (2015).
+We analyze both the prediction outcomes and error disparity in age and openness prediction for both models: HaRT, which considers individual context, and GRIT, which incorporates both individual and group context.
+First, we split the task test dataset into different buckets based on the age groups (specifically, <18, 18-21, 21-30, 30-45, and >45 years old) of the users in the test set, and then we compare the performance of our models across these buckets.
+Results from Table 4 indicate that pre-training with individual and group contexts together performs better for estimating age across all the age groups, which implies it makes fewer errors as a function of the socio-demographic attribute age.
+We see similar trends for assessing openness personality (see Appendix Tables 6 and 8), suggesting that the group attribute prediction may act as a regularizer for models pre-trained with both individual and group contexts, thus aiding the models to make fewer errors across all age buckets.
+To further confirm, we compute the mean error disparity ($MED$) as the sum of the differences in the performance metric (Pearson correlation for age, and disattenuated Pearson correlation for openness) across each pair of age buckets, which is then averaged by the number of pairs (Shah et al., 2020).
+A lower averaged sum of differences implies fewer errors as a function of the age groups.
+Lower $MED$ scores for models pre-trained with individual and group context in Table 5 support our previous error analysis.
 7 Conclusion
-NLP benefits from modeling latent human context, such as sociodemographic group traits or individual tendencies.
-A recent development has been to incorporate this additional human context into the PLMs’ pre-training regimen.
+NLP benefits from modeling latent human context, such as socio-demographic group features or individual traits.
+A recent development has been to incorporate this additional human context into the pre-training regimen of LMs.
 However, humans exhibit varying degrees of group and individual characteristics.
-Learning about the impacts of pre-training with different types of human context will propel us forward in including human context in our base LLMs (Soni et al., 2023).
-To assess its impacts, we compare three types of PLMs pre-trained with sociodemographic group attributes, individual human contexts, or a combination across five user- and document-level tasks.
-We find that pre-training with individual and group human context improves the two user-level tasks: age and personality prediction.
-Pre-training with individual human context improves the three single-document classification tasks, including stance and topic detection.
-Pre-training solely on group context helps group-based document classification tasks, though not optimally.
-Our findings show a promising step towards modeling human context and provide valuable insights for the NLP community to think of additional strategies for improving models with task-dependent human context in pre-training.
+Understanding the impacts of pre-training with different types of human context will advance the integration of human context into our base LLMs (Soni et al., 2024).
+To assess the impacts, we compare three types of PLMs pre-trained with socio-demographic group attributes, individual human contexts, and combined group and individual traits, across five user- and document-level tasks.
+Our findings indicate that pre-training with both individual and group human context improves the two user-level regression tasks: age and personality prediction.
+Pre-training with individual human context enhances the performance of the three single-document classification tasks, including stance and topic detection.
+Interestingly, inclusion of both individual and group attributes results in reduced performance on the text classification tasks.
+Meanwhile, pre-training solely on group context aids in group-based document classification tasks, albeit suboptimally.
+These results represent a promising step towards modeling human context and offer valuable insights for the NLP community to investigate additional strategies for improving models with task-dependent human context during pre-training.
 Limitations
-The purpose of our study is to compare the impacts of modeling sociodemographic group attributes and modeling individual user traits, and we use relevant models to represent each of the approaches.
+The purpose of our study is to compare the impacts of modeling socio-demographic group attributes and modeling individual user traits, and we use relevant models to represent each of the approaches.
 There are likely to be other ways to model these approaches and the models we use are only one of the ways.
@@
 Ethical Considerations
-Models that incorporate sociodemographic information need to be considered with special scrutiny.
+Models that incorporate socio-demographic information need to be considered with special scrutiny.
 On the one hand, they have the potential to produce fairer and more inclusive results, because they can account for human language variation.
@@
 Acknowledgments
-This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #585968.
-The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S.
+This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005, and a grant from the CDC/NIOSH (U01 OH012476).
+The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, any other government organization, or the U.S.
 Government.
@@
 Appendix A Appendix
-A.1 Experimental Settings
-We closely follow the experimental settings from Soni et al.
-(2022) and similarly use Optuna framework (Akiba et al., 2019) for hyperparameter search.
-We search for learning rates between 5e-6 and 5e-4, and between 1e-7 and 1e-5 for different tasks.
-We will make our best found hyperparameter values publicly available with our code and results in the github repository.
-All experiments are run on NVIDIA RTX A6000 GPUs of 48GB.
-Pre-training takes approx 14 hours for 1 epoch and fine-tuning takes approx 1-4 hours depending on the task.
+A.1 Pre-training GRIT
+Pre-training data.
+We use a subset of the pre-training data for HaRT, consisting of the demographics and personality information.
+This subset contains the Facebook posts from Park et al.
+(2015) as used by Soni et al..
+Our dataset is consistent with the inclusion criteria for HaRT to ensure moderate language history for each user: we include English posts from users with at least 50 total posts and at least 1000 words.
+This dataset consists of just over 63,000 unique users, which we split into a training dataset consisting of messages from 56,930 users, a development dataset that consists of messages from 1836 users that were not part of the training set, and a test set of messages from a separate set of 4438 users that are neither in training nor the development set.
+To evaluate the human attribute prediction in GRITope, we use a subset of the test set consisting of messages from 1745 users to accommodate for questionnaire reliability.
+We use the Facebook posts for the HuLM task and the demographic and personality scores of consenting Facebook users (Kosinski et al., 2013) for the human attribute prediction task.
 Age bucket #Users HaRT (Ind) GRITage (Ind+Grp) GRITope (Ind+Grp)
@@
 Table 8: Mean squared error for openness over five age buckets using different types of human contexts for error analysis. Bold indicates best in row (lower error is better).
+Training.
+We use HaRT’s pre-trained weights as the base weights for GRIT and randomly initialize the newly introduced weights for human attribute prediction.
+GRIT is trained on our pre-training dataset using the 5e-5 learning rate after experimenting with a few learning rates, including that used for HaRT’s pre-training.
+Following HaRT, and due to computing limitations, each training instance is capped to 8 blocks of 1024 tokens each, with train batch size as 1 per device and evaluation batch size as 20 per device, trained over 2 GPUs for eight epochs.
+We explored multiple joint losses before resorting to the homoscedastic loss computation.
+Since HaRT caps to 4 train blocks for user-level downstream tasks, we also pre-train GRITage and GRITope with four training blocks.
+A.2 Experimental Settings
+We closely follow the experimental settings from Soni et al.
+(2022) and similarly use Optuna framework (Akiba et al., 2019) for hyperparameter search.
+We search for learning rates between 5e-6 and 5e-4, and between 1e-7 and 1e-5 for different tasks.
+We will make our best found hyperparameter values publicly available with our code and results in the github repository.
+All experiments are run on NVIDIA RTX A6000 GPUs of 48GB.
+Pre-training takes approx 14 hours for 1 epoch and fine-tuning takes approx 1-4 hours depending on the task.